### v3

# A General Framework for Joint Multi-State Models

Abstract
Conventional joint modeling approaches generally characterize the relationship between longitudinal biomarkers and discrete event occurrences within terminal, recurring or competing risk settings, thereby offering a limited representation of complex, multi-state trajectories.
We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs. The proposed framework also accomodates nonlinear longitudinal submodels and scalable inference via stochastic gradient descent. This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented. The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
We derive the complete likelihood, model selection criteria, and develop scalable inference procedures based on stochastic gradient descent to enable high-dimensional and large-scale applications. In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
Through simulation and application to the PAQUID cohort, we demonstrate accurate parameter recovery and individualized prediction.
keywords:
joint modeling, multi-state processes, longitudinal data, survival analysis, stochastic gradient descent, dynamic prediction
1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [papageorgiou2019overview], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [rizopoulos2012joint], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017]. To address these limitations, specialized frameworks have been developed to handle competing risks and recurrent events. However, these remain constrained to specific event types and fail to accommodate more general multi-state transitions. In such settings, multi-state models provide a more general and flexible approach [ferrer2016joint].
Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [lovblom2024modeling]. In biostatistics, these models are widely used for survival and reliability analysis, allowing for a richer and more accurate representation by capturing alternative paths to an event of interest, intermediate events, and progressive disease. Key components of multi-state models include transition intensity functions, which denote the instantaneous risk of moving from one state to another, and transition probability functions, which describe the probability of transition over longer intervals. Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [asanjarani2022estimation]. This assumption can be relaxed by adopting a semi-Markov formulation, in which the transition probabilities depend not only on the current state but also on the sojourn time, effectively resetting the time scale after each transition.
The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework . While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with an acyclic time-to-event process.
The model proposed by Ferrer et al. [ferrer2016joint] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen (PSA) measurements) and a multi-state process (e.g., clinical progressions in prostate cancer). These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between the continuous longitudinal biomarker trajectory and the discrete transitions between health states.
The model we propose operates on an arbitrary directed graph, enabling representation of complex state transitions and recurrent events. We derive the complete likelihood for the nonlinear joint model, introduce an efficient stochastic approximation inference method, and develop dynamic prediction tools. To illustrate its practical relevance, we conduct both a simulation study and a synthetic biomedical case study.
This paper is organized as follows. Section 2 reviews the background on multi-state and joint modeling frameworks. Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and proposed model selection creteria. Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework. Section 6 reports simulation experiments that assess parameter recovery and convergence properties of the proposed model, whereas Section 7 applies the framework to data from the PAQUID cohort. Finally, Section 8 concludes with perspectives and future research directions.
2 Background
Joint modeling of longitudinal and time-to-event data provides a unified statistical framework to analyze the interplay between continuous biomarker dynamics and event occurrence processes. In its classical form, this framework couples a mixed-effects model describing individual biomarker trajectories with a survival model for event times [rizopoulos2012joint]. Such models have become central in biomedical research, particularly for dynamic prediction and personalized risk assessment [papageorgiou2019overview]. However, these traditional joint models generally focus either on a single terminal event, competing, or recurrent events, limiting their ability to represent more complex event histories that include intermediate, recurrent, or competing events. To address these limitations, multi-state extensions have been developed, providing a natural framework for describing transitions between multiple clinical or biological states over time [ferrer2016joint, KrolMauguen2017].
2.1 Multi-State Markov Processes
A multi-state stochastic process is defined as a process $S(t)$ for $t\geq 0$, where $S(t)$ can take a finite number of values (states), often labeled $1,2,\dots,p$. Quantities of interest typically include the probability of being in a certain state at a given time and the distribution of first passage times.
A Markov process (or continuous-time Markov chain) is a specific class of multi-state models in which future transitions between states depend only upon the current state. This Markov property means the process is memoryless. A key consequence is that the duration spent in any state follows an exponential distribution, implying a constant hazard rate for leaving that state [jackson2011multi, putter2007competing].
However, the Markov assumption can be unrealistic in many real-world applications. For example, in the study of human sleep stages, sojourn times often do not follow an exponential distribution [dong2008non, roever2010statistical], and in the case of chronic diseases such as AIDS, the risk of disease progression can depend on the time elapsed since infection [andersen1999multi]. To address these limitations, semi-Markov processes (SMPs) were introduced, which allow for arbitrary distributions of sojourn times while retaining the Markov property for the embedded discrete-time chain. This flexibility makes SMPs suitable for modeling complex disease progression and patient recovery scenarios [putter2007competing].
2.2 Multi-State Semi-Markov Processes
Multi-state semi-Markov processes (MSMPs) offer a natural generalization by allowing the distribution of sojourn times to be arbitrary.
In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions. The transition intensities $\lambda_{k^{\prime}\mid k}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$ are therefore assumed to be non-identically zero. The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
MSMPs have been widely applied in a range of disciplines. In reliability, they are used to modeling degradation and repair processes [limnios2001semi]. In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [commenges2006likelihood, fiocco2008modelling]. Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [luciano2006semi]. In these contexts, MSMPs retain the Markov property in the embedded jump chain while offering increased realism through flexible dwell time modeling.
From a methodological standpoint, MSMPs extend estimation strategies developed for Markov models, such as maximum likelihood or Bayesian inference, and can accommodate interval-censored or misclassified data [frydman2005estimation]. This makes them a powerful and general tool for multi-state event history analysis.
While both joint models and multi-state models offer powerful tools, they primarily address different facets of disease progression and have complementary strengths. Multi-state models excel at understanding the sequence and timing of discrete events, while joint models are adept at modeling continuous biomarker trajectories and their associations with event outcomes for dynamic prediction. The complexity of real-world diseases often necessitates a more integrated approach that can leverage the strengths of both frameworks. For instance, in prostate cancer, multiple types of relapse may occur successively (a multi-state process), and their risk is influenced by the dynamics of longitudinal biomarkers like PSA.
2.3 Joint Models and Multi-State Processes
Previous research has initiated such integration, with models like the one proposed by Ferrer et al. [ferrer2016joint], which combines a linear mixed sub-model for longitudinal data with a multi-state sub-model using shared random effects. This model enables the simultaneous analysis of repeated measurements of a biomarker (the longitudinal process, e.g., PSA levels) and the times of transitions between multiple health states (the multi-state process). It explicitly accounts for the link between these two correlated processes and uses information from the biomarker dynamics to explain or predict clinical progression events. The multi-state sub-model assumes a non-homogeneous Markov process for the clinical progression, meaning that the future of the process depends only on the current state (Markov property) and the time elapsed since study entry (non-homogeneous property).
Although Ferrer et al. [ferrer2016joint] note that their joint modeling framework could be adapted to semi-Markov processes in other contexts, their primary application to prostate cancer progression assumes a non-homogeneous Markov multi-state process. However, this framework is not well suited for arbitrary transition patterns other than the ones modeled by DAGs.
While both joint models and multi-state models offer powerful tools for analyzing disease progression, they emphasize different aspects of the underlying process. Multi-state models capture the sequence and timing of discrete transitions between states, whereas joint models focus on the evolution of continuous longitudinal markers and their association with event risk. Existing joint multi-state models, such as that of Ferrer et al. [ferrer2016joint], are restricted to linear mixed-effects submodels and to acyclic Markov transition graphs, as their likelihood relies on a sequential factorization that is not valid when transitions form cycles. However, many biological processes involve cyclic or semi-Markov transitions and nonlinear biomarker dynamics. These cannot be handled by existing frameworks, motivating the present general formulation.
In contrast, many real-world processes—such as recurrent disease relapse or recovery—require more flexible representations involving nonlinear biomarker dynamics, semi-Markovian dependencies, and recurrent cycles. The next section introduces a general framework that integrates these features within a unified likelihood-based formulation, enabling efficient inference and dynamic prediction along complex event trajectories.
3 A General Multi-State Joint Modeling Framework
To address the limitations imposed by traditional joint modeling frameworks, we extend them to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t\geq 0}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process. Both components are linked through shared latent structures, and the event process is represented on an arbitrary directed graph supporting both Markovian and semi-Markovian transition mechanisms.
3.1 Notation Conventions
Throughout this section, $i=1,\dots,n$ indexes individuals, and $t_{ij}$ denotes the $j$-th observation time of individual $i$. The vector $Y_{ij}\in\mathbb{R}^{d}$ represents longitudinal biomarker measurements, and $X_{i}\in\mathbb{R}^{p}$ the associated covariates.
The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, while $\psi_{i}\coloneqq f(\gamma,X_{i},b_{i})$ denotes the subject-specific parameters in the longitudinal submodel. The filtration $\mathcal{H}_{i}(t)$ represents the observed history of the longitudinal process up to time $t$.
Transition times and states are denoted by $(T_{i\ell},S_{i\ell})$, with right-censoring time $C_{i}$. We also define the index of the last observed transition before time $t$ denoted by $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$.
3.2 Extending Multi-State Joint Models
The joint model proposed by Ferrer et al. [ferrer2016joint] established a foundational link between longitudinal biomarkers and multi-state processes through shared random effects. However, their formalism relies on binary censoring indicators to characterize transitions, thereby inherently restricting the modeling framework to Directed Acyclic Graphs (DAGs) and excluding recurrent events. To overcome this limitation, we present a new method based on arbitrary graph transitions. By shifting away from binary indicators toward a more flexible graph-theoretic specification, our approach accommodates complex disease trajectories, including recurrent events and cyclic transitions that cannot be captured within a DAG-based structure.
3.2.1 Graph Structure
Let $G=(V,E)$ be a directed graph, where $V=\{1,\dots,p\}$ denotes the set of states and $E\subset V\times V$ the set of admissible transitions. The graph encodes all possible paths of the multi-state process, allowing for competing, recurrent, or absorbing transitions. Figure 1 illustrates an example of a four-state model including an absorbing state (Death). The corresponding adjacency matrix $A=(A_{kk^{\prime}})_{1\leq k,k^{\prime}\leq|V|}$ satisfies $A_{kk^{\prime}}=1$ if $(k,k^{\prime})\in E$ and 0 otherwise.
Figure 1: Example of a 4-state transition graph $G=(V,E)$, including an absorbing state (4).
Such a representation generalizes standard joint models, which can be recovered as special cases: single-event (linear chain), competing-risks (two absorbing transitions), or recurrent-event settings (cyclic transitions) as illustrated in Figure 2.
Standard Risk
Competing Risks
Recurrent Risk
Figure 2: Illustration of classical joint modeling approaches: (left) standard risk, (middle) competing risks, and (right) recurrent risk.
Such joint modeling offers a flexible framework that extends beyond classical approaches and is particularly useful in medical applications.
3.2.2 Individual Trajectories
Each individual $i$ follows a latent trajectory
$\mathcal{T}_{i}^{*}\coloneqq\big((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\big),$
where $T_{i\ell}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state. The observed trajectory is right-censored at time $C_{i}$, so that only $\{(T_{i\ell},S_{i\ell}):T_{i\ell}\leq C_{i}\}$ is observed. Between transitions, the state $S_{i}(t)$ is piecewise constant. We note $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$ the index of the last observed transition before time $t$ and $m_{i}=m_{i}(C_{i})$.
The process is assumed to satisfy a (possibly time-inhomogeneous) semi-Markov property
$\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid\{(T_{i\ell^{\prime}},S_{i\ell^{\prime}})\}_{\ell^{\prime}\leq\ell}\big)=\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid(T_{i\ell},S_{i\ell})\big),$ (1)
which is relaxed compared to a stricter Markov property by allowing sojourn-time dependence.
3.2.3 Transition Intensities and Survival Functions
For any admissible transition $(k,k^{\prime})\in E$, let $\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})$ denote the instantaneous risk of moving from $k$ to $k^{\prime}$ at time $t$, given entry time $t_{0}$
$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
The cumulative risk and the corresponding survival function are
$\Lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\int_{t_{0}}^{t}\lambda_{i}^{k^{\prime}\mid k}(w\mid t_{0})\,dw,$
$\mathbb{P}(T_{i,\ell+1}>t\mid T_{i\ell},S_{i\ell})=\exp\left[-\sum_{s:(S_{i\ell},s)\in E}\Lambda_{i}^{s|S_{i\ell}}(t\mid T_{i\ell})\right].$
Conditional transition probabilities are
$\mathbb{P}(S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1},T_{i\ell},S_{i\ell})=\frac{\lambda_{i}^{k^{\prime}|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}{\sum_{s:(S_{i\ell},s)\in E}\lambda_{i}^{s|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}.$
3.3 Multi-State Joint Model Specification
The proposed multi-state joint model with Gaussian prior and homoscedastic Gaussian noise is specified by:
$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q),\\ \lambda_{i}^{k^{\prime}\mid k}\left(t\mid X_{i},T_{im_{i}(t)},\mathcal{H}_{i}(t)\right)=\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})\exp\big(\alpha^{k^{\prime}\mid k}g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})+\beta^{k^{\prime}\mid k}X_{i}\big).\end{cases}$
Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k^{\prime}\mid k}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k^{\prime}\mid k}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$. This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications.
The coefficient $\alpha^{k^{\prime}\mid k}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$. More concretely, for a one-unit increase in $g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k^{\prime}\mid k})$. Likewise, $\beta^{k^{\prime}\mid k}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k^{\prime}\mid k}_{j})$. Thus, $\alpha^{k^{\prime}\mid k}$ captures the degree to which the biomarker trajectory (via its latent parameter $\psi_{i}$ or function $g^{k^{\prime}\mid k}$) influences the transition risk, while $\beta^{k^{\prime}\mid k}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.
3.3.1 Model Variants
Two baseline hazard conventions are common:
$\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})=\begin{cases}\lambda_{0}^{k^{\prime}\mid k}(t-T_{im_{i}(t)})&\text{(clock reset)},\\ \lambda_{0}^{k^{\prime}\mid k}(t)&\text{(clock forward)}.\end{cases}$
The clock-reset form models risk as a function of time spent in the current state, whereas the clock-forward form measures risk with respect to global time since study entry.
3.3.2 Likelihood Formulation
To derive the marginal likelihood, we decompose the joint distribution of the longitudinal and multi-state processes under a set of standard conditional independence assumptions detailed in Table 3.3.2 below.
Assumptions for Likelihood Factorization A. Latent-level independence A1. Random effects $(b_{i})_{i}$ are mutually independent across individuals. B. Conditional independence within individuals B1. Longitudinal observations $(Y_{ij})_{ij}$ are mutually independent given $b_{i},X_{i}$. B2. Trajectories $(\mathcal{T}_{i}^{*})_{i}$ are mutually independent given $b_{i},X_{i}$. B3. Conditional on $b_{i},X_{i}$, the longitudinal and event processes are mutually independent. C. Censoring and process assumptions C1. Censoring times $(C_{i})_{i}$ are mutually independent and noninformative, i.e. $C_{i}\mathrel{\perp\!\!\!\perp}(Y_{i},\mathcal{T}_{i}^{*})\mid b_{i},X_{i}$. C2. Conditionally on $b_{i},X_{i}$, event trajectories satisfy the semi-Markov property 1: $p(\mathcal{T}_{i}^{*}\mid b_{i},X_{i})=\prod_{\ell\geq 0}p\left((T_{i,\ell+1},S_{i,\ell+1})\mid T_{i\ell},S_{i\ell},b_{i},X_{i}\right)$.
These grouped assumptions mirror those used in classical joint modeling [rizopoulos2012joint] and multi-state survival analysis [putter2007competing]: independence across subjects; the conditional independence structure linking the longitudinal and event submodels through shared random effects; noninformative censoring and (semi-)Markovian dynamics.
Let $\theta=(\gamma,Q,R,\alpha,\beta)$ denote the vector of model parameters. For subject $i$, we observe the longitudinal measurements $Y_{i}=(Y_{i1},\dots,Y_{in_{i}})$ and the event trajectory $\mathcal{T}_{i}=\{(T_{i\ell},S_{i\ell})_{0\leq\ell\leq m_{i}}\}$. The joint likelihood then factorizes as
$p(Y_{i},\mathcal{T}_{i}\mid X_{i},C_{i},\theta)=\int p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta)\,db_{i},$
where $b_{i}$ are individual random effects. This formulation generalizes the joint likelihoods of wulfsohn1997joint and ferrer2016joint to arbitrary multi-state event structures. Each part can then be explicitely expressed using Assumptions 3.3.2, yielding an expression very similar to that obtained by rizopoulos2012joint.
Prior likelihood:
$p(b_{i}\mid\theta)=\frac{(2\pi)^{-q/2}}{\det(Q)^{1/2}}\exp\left(-\tfrac{1}{2}b_{i}^{T}Q^{-1}b_{i}\right),$
Longitudinal likelihood:
$p(Y_{i}\mid b_{i},X_{i},\theta)=\prod_{j=1}^{n_{i}}\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
$\text{with }\psi_{i}=f(\gamma,X_{i},b_{i}).$
Semi-Markov likelihood:
$p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)=$ $\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid b_{i},X_{i},(T_{il},S_{il}),\theta\right)$ (2)
$\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{k^{\prime}\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right).$
The proof of the expression of the Semi-Markov likelihood (2) is provided in Appendix B.
The likelihood factorization above is valid and forms the basis for scalable inference procedures described in Section 4, leveraging the complete trajectory of biomarkers to refine predictions of transitions and survival [de2010mstate].
Moreover, here, we assume that the initial state $S_{i0}$ is observed. However, the framework could also incorporate a multinomial model for unobserved initial states [yiu2018clustered].
3.4 Model Selection and Information Criteria
To compare competing model specifications, we rely on information criteria based on the marginal likelihood as defined in Section 4. The Akaike Information Criterion (AIC) [akaike2003new] and the Bayesian Information Criterion (BIC) [schwarz1978estimating] are standard tools. While the AIC relies on an asymptotically unbiased estimator of the $\log$-likelihood, the BIC aims to identify the true model with great probability as the sample size grows.
The Akaike Information Criterion is computed as follows
$\mathrm{AIC}\coloneqq-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
where $k$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
The Bayesian Information Criterion is similarly defined but imposes a stronger penalty for model complexity, growing in $\log n$ where $n$ is the number of observations. Particularly in the context of mixed effects models, and even more so for multi-state joint models, the number of observations $n$ may vary from one definition to another, either the total number of repeated measurements or the number of individuals [delattre2014note]. To alleviate this problem, we recall the derivation of the BIC from the Laplace approximation [lebarbier2006introduction]. For a prior distribution $\pi$ on a set of models $\mathcal{M}$, the Laplace approximation approaches up to constant terms which could also be accounted to improve the approximation, the posterior probability given observed data $x$ defined in Section 4 by
[…]
where $H_{\hat{\theta}}$ denotes the Hessian matrix of the marginal $\log$-likelihood evaluated at the Maximum Likelihood Estimator (MLE). Then, under standard regularity conditions that allow differentiation and integration to be interchanged twice, we have that
[…]
Thus, our BIC criterion can be written as
[…]
In both cases, one should aim at minimizing the chosen criterion.
4 Statistical Inference
The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\text{complete}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\text{marginal}}$, obtained by integrating them out
$\mathcal{L}_{\text{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)\,db,$
where
$\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta).$
It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model, necessitating numerical approximation methods such as Monte Carlo integration or Laplace approximation.
Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace] and Gauss–Hermite quadrature. These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic], alongside with MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo.
Under interchangeability of integral and differentiation, setting $x\coloneqq(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
$\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta;x,b)\right).$
This expectation is approximated using Monte Carlo samples from the posterior $p(\cdot\mid x,\theta)$, avoiding the need to evaluate the intractable marginal likelihood $\mathcal{L}_{\text{marginal}}(\theta;x)$. This formulation naturally supports stochastic gradient ascent with MCMC-based posterior sampling and is amenable to minibatch parallelization across individuals, ensuring the convergence to a critical point. The update rule is as follows:
Algorithm 1 Stochastic gradient ascent model inference
1:data $x$; initial parameter $\theta^{(0)}$; step sizes $(\eta_{t})_{t\geq 0}$ with $\sum_{t}\eta_{t}=+\infty$, $\sum_{t}\eta_{t}^{2}<+\infty$; batch size $K$; sampler $p_{t}^{\mathrm{MCMC}}$; stopping criterion.
2:
3:while not converged do
4:  Draw $b_{k}\sim p(\cdot\mid x,\theta_{n})\approx p_{n}^{\text{MCMC}}$.
5:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\frac{\eta_{t}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta^{(t)};x,b_{k})$.
6:end while
7:return $\hat{\theta}$
Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient descent framework, or even specialized optimizers such as Adam [kingma2014adam], NAdam [dozat2016incorporating] or Adagrad [duchi2011adaptive]. In particular, the Adam optimizer will later be used in our numerical experiments.
5 Dynamic prediction
The simulation of individual trajectories using Algorithm 2 enables prediction in the multi-state joint modeling framework. Specifically, predictions are derived from simulated event trajectories, allowing for individualized risk assessment and forecasting of future states or event times based on a subject’s observed biomarker and event history. This approach generalizes classical dynamic prediction by leveraging the rich structure of multi-state processes and the joint distribution of longitudinal and event data.
Suppose we are interested in some functional of the true (unobserved) future trajectory,
$\chi^{*}(\mathcal{T}_{i}^{*})\in\mathbb{Y},$
where $\mathbb{Y}$ is some generic space.
For example, one can think of the state taken by the individual $i$ at a time $u$. As is the case in traditional dynamic prediction however, we are only interested in predicting quantities or characteristics that depend on a yet unobserved future, given prior information, i.e. the trajectory up to some prediction time $t\leq u$ and observed longitudinal markers. Contrary to (single transition) joint models, the conditional probability distribution on $(\mathbb{R}\times\mathcal{S})^{\mathbb{N}}$ cannot be analytically derived.
Nonetheless, as seen in Algorithm 2, we are able to accurately simulate each trajectory. Therefore, a Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its expected value:
$\hat{\chi}_{i}=\coloneqq\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
In general however, $\chi^{*}$ may very well depend on (countably) infinitely many transitions. As a result, since the proposed estiamtion relies on simulted samples, we require the quantity to depend only on a finite subset of transitions.
Assumption 1.
Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}\coloneqq\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
$\begin{cases}\tau_{i}<+\infty\text{ a.s.},\\ \chi^{*}(\mathcal{T}_{i}^{*})=\chi\left((T_{il},S_{il})_{1\leq l\leq\tau_{i}}\right)\end{cases}$
Essentially, given Assumption 1, with probability one we are guaranteed that the quantity of interested may be computed in a finite number of simulation steps. Indeed, the prediction algorithm for a single sample may be summarized as follows:
[…]
2:
[…]
4:while $n<\tau_{i}$ do
5:  Append simulated $(T_{in},S_{in})$ to $\mathcal{T}_{i}$ if $n>\operatorname{len}(\mathcal{T}_{i})$
6:  $n\leftarrow n+1$
7:end while
8:return $\chi(\mathcal{T}_{i})$
Numerous quantities of interest may be encompassed by this framework. We give multiple examples below. First, for a directed acyclic graph $G=(V,E)$, we define the depth of $G$, denoted by $\operatorname{depth}(G)$, as the length of the longest directed path in $G$, that is,
$\operatorname{depth}(G)\coloneqq\max_{(v_{0},v_{1},\dots,v_{k})}k,$
where $(v_{0},v_{1},\dots,v_{k})$ ranges over all directed paths in $G$.
Example 1 (State at time $u$).
Let $G=(V,E)$ be a directed acyclic graph, $u\in\mathbb{R}$ be a fixed time, and $\mathbb{Y}=V$ be the set of possible states. Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq u\text{ or }S_{in}\text{ is absorbing}\}$. Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$. Then, take $\chi_{u}^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$, which corresponds to the state of the individual $i$ at time $u$. In particular, if $V=\{0,1\}$ and $E=\{(0,1)\}$, we recover the special case used for survival probability estimation in standard joint models [rizopoulos2011dynamic].
Example 2 (Hitting time).
Let $G=(V,E)$ be a directed acyclic graph, $\mathbb{A}\subset V$ a non-empty subset of states, and $\mathbb{Y}=\mathbb{R}\cup\{+\infty\}$. For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subset V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$. Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\vee\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$. Then, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$ and take $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})=T_{i\tau_{i}}+\mathds{1}_{\{S_{in}\}\not\rightsquigarrow\mathbb{A}}(+\infty)$ represents the hitting time for the set $\mathbb{A}$.
In Example  1, the stopping time $\tau_{i}$ captures the step at which the process for individual $i$ reaches or exceeds a given time $u$, or enters an absorbing state. The resulting value $\chi^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$ therefore represents the state of the individual at that specific time.
In Example 2, the stopping time $\tau_{i}$ corresponds to the first time the process reaches a target subset of states $\mathbb{A}$, or becomes unable to reach it in the future. The associated value $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})$ thus represents the hitting time of the set $\mathbb{A}$, which may be finite if $\mathbb{A}$ is reached, or $\infty$ otherwise.
Other practical applications may include finding the expected number of edges in some given trajectory, the number of times an individual has returned in a specific state, the expected time between transitions…
This simulation-based dynamic prediction framework generalizes classical approaches from single-event joint models to the multi-state context. It enables individualized forecasting of future state occupancy, event risks, sojourn times, and other clinically relevant outcomes, fully exploiting the subject’s observed biomarker trajectory and event history. In Section 7, we demonstrate its application to forecasting dependency trajectories in the PAQUID cohort.
6 Simulation Study
The purpose of this simulation study is to evaluate the finite-sample performance and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known. We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model. This setup allows us to assess both statistical accuracy and computational scalability of the stochastic-gradient estimation procedure.
All simulations, estimations, and figures reported in this section were produced using the open-source jmstate Python package, available on PyPI, which we developed to implement the multi-state joint modeling framework and inference algorithms described in Sections 3.2 and 4.
6.1 Model Specification
The simulated model involves three states—Healthy (0), Sick ($1$), and Terminal (22)—connected by the transitions $0\rightarrow 1$, $0\rightarrow 2$, and $1\rightarrow 2$ (Figure 3). This configuration captures the typical monotone evolution of a disease process, with both direct and indirect paths to the terminal state. Transition intensities follow an exponential baseline hazard under a clock-reset convention.
The longitudinal process follows a continuous piecewise-affine regression:
$h(t,\psi)=\psi_{1}+\psi_{2}t+\mathds{1}_{t>\tau}(\psi_{3}-\psi_{2})(t-\tau),$
with a known inflection time $\tau=6$. The individual-specific parameters $\psi_{i}=(\psi_{i1},\psi_{i2},\psi_{i3})$ arise from a nonlinear mixed-effects model with Gaussian random effects $b_{i}\sim\mathcal{N}(0,Q)$ and independent Gaussian noise $\epsilon_{ij}\sim\mathcal{N}(0,R)$. The link function combines the current biomarker value and its slope,
$g(t,X_{i},\psi_{i})=\big(h(t,\psi_{i}),\frac{\partial h}{\partial t}(t,\psi_{i})\big),$
shared across transitions. Covariates $X_{i}$ are normally distributed and enter linearly through $\beta^{k^{\prime}\mid k}X_{i}$.
Longitudinal data were collected at up to $20$ random time points per subject and censored at $C_{i}\sim\mathcal{U}([10,15])$. Censoring also truncates the biomarker trajectories. The true parameter vector lies in $\mathbb{R}^{16}$. Each simulated dataset comprised $n=1000$ individuals, generating approximately balanced transition counts (Figure 4).
This trajectory mimics a biomarker whose evolution slows after treatment or onset of symptoms, while preserving continuity at the change point.
Figure 3: State transition diagram of the simulated three-state model.
6.2 Estimator Convergence and Accuracy
Parameter estimation was performed using the stochastic-gradient inference procedure of Section 4, with the Adam optimizer (learning rate $0.5$, batch size $K=100$) and an adaptive stopping criterion based on the first and second moments of parameter differences. Convergence was typically achieved in 400400500500 iterations and required around $12$ seconds per run on an AMD Ryzen 5 processor.
Across $n_{\text{runs}}=100$ independent replications, the mean relative bias of all parameters remained below $2\%$, and the root-mean-square error (RMSE) scaled approximately as $n^{-1/2}$. Parameters governing the biomarker trajectory $(\gamma,Q,R)$ were consistently well identified, while association coefficients $(\alpha^{k^{\prime}\mid k},\beta^{k^{\prime}\mid k})$ exhibited slightly higher variability for rare transitions. Representative convergence trajectories are shown in Figure 10, and quantitative results are summarized in Table 5.
Furthermore, longitudinal measurements were taken at $m_{i}\leq 20$ time points, and censoring times were drawn from $C_{i}\sim\mathcal{U}\left([10,15]\right)$. Longitudinal observations were also truncated at the censoring times.
A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below.
From $\rightarrow$ To # Observations
$0\rightarrow 1$ 613
$0\rightarrow 2$ 375
$1\rightarrow 2$ 592
Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions between states for the complete population of $1000$ individuals.
6.3 Estimator Convergence
To illustrate the convergence of the estimator, we examine the optimization process for a particular run with $n=1000$ (see Figure 10), as well as RMSE values computed from $n_{\text{runs}}=100$ independent runs on different simulated datasets, all generated from the same underlying distribution. The results of these simulations are shown in Table 5.
To fit the model, we first have to define initial parameters, which can be particularly important for the optimization process if the model is not well-behaved. In the present case, without prior knowledge of the true parameters, we zero the values and use the identity matrix for both covariance matrices.
Formally:
$m_{i}^{(t)}\leftarrow\beta_{i}m_{i}^{(t-1)}+(1-\beta_{i})(\theta^{(t)}-\theta^{(t-1)})^{i},\quad\hat{m}_{i}^{(t)}=\frac{m_{i}^{(t)}}{1-\beta_{i}^{t}},$
and the optimization process is terminated when
$|\hat{m}_{1}^{(t)}|\leq 10^{-6}+10^{-1}\sqrt{\hat{m}_{2}^{(t)}}.$
The Figure 10 shows the evolution of the parameters during the optimization process, and Table 5 summarizes the accuracy of the inferred parameters.
Parameter True value Mean Standard error RMSE
$\gamma_{1}$ 2.5000 2.4960 0.0370 0.0372
$\gamma_{2}$ -1.3000 -1.3039 0.0257 0.0260
$\gamma_{3}$ 0.2000 0.1936 0.0267 0.0275
$\tilde{Q}_{1}$ 0.2554 0.2243 0.0442 0.0540
$\tilde{Q}_{2}$ 0.8047 0.8009 0.0260 0.0263
$\tilde{Q}_{3}$ 0.6020 0.6001 0.0276 0.0277
$\tilde{R} -$0.2653 -0.2684 0.0083 0.0089
$\alpha_{1}^{1\mid 0}$ -0.5000 -0.49997 0.0486 0.0486
$\alpha_{2}^{1\mid 0}$ -3.0000 -2.9962 0.0800 0.0801
$\alpha_{1}^{2\mid 0}$ -1.0000 -0.9990 0.0847 0.0847
$\alpha_{2}^{2\mid 0}$ -5.0000 -4.9897 0.1191 0.1196
$\alpha_{1}^{2\mid 1}$ 0.0000 0.0011 0.0327 0.0327
$\alpha_{2}^{2\mid 1}$ -1.2000 -1.2045 0.0429 0.0432
$\beta^{1\mid 0}$ -1.3000 -1.3015 0.0501 0.0501
$\beta^{2\mid 0}$ -0.9000 -0.8982 0.0670 0.0670
$\beta^{2\mid 1}$ -0.7000 -0.6981 0.0551 0.0551
Figure 5: Comparison of true and estimated parameters (averaged over 100 runs).
The simulation confirms that the proposed inference method accurately recovers all parameters of the multi-state joint model under realistic sample sizes. The optimization demonstrates stable convergence (Figure 10), with minimal sensitivity to initialization. The algorithm’s computational efficiency allows for extensive simulation or cross-validation studies at negligible cost.
7 Application to the PAQUID Cohort
7.1 Introduction
The data used in this section originates from the PAQUID cohort [Letenneur1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging. A subsample of $n=500$ individuals was followed over a period of up to $20$ years [Proust-Lima2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics. In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE, see Figure 6), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency). In this work, we focus on the association between cognitive decline and the progression of physical dependency by jointly modeling the longitudinal trajectory of MMSE scores and the transitions between the four HIER states using a joint multi-state model. More specifically, the state of an individual $i$ at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$. This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
Both time and longitudinal values were normalized in [1,1][-1,1]. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
From $\rightarrow$ To # Observations
$0\rightarrow 1$ 85
$1\rightarrow 2$ 135
$2\rightarrow 3$ 86
$0\rightarrow 2$ 44
$1\rightarrow 3$ 15
$0\rightarrow 3$ 11
Figure 6: On the left, each light gray line represents one of the $500$ observed sequences of (normalized) MMSE scores, and three functional clusters are also represented; on the right, observed transitions between states.
Therefore, the chosen transition graph is as follows, and only allows monotonic transitions from a lower level of physical dependency to a higher one.
Figure 7: Schematic representation of the four ordered HIER states, with possible transitions. HIER quantifies physical dependency (0: no dependency to 33: severe dependency) [Proust-Lima2017].
We consider an exponential baseline hazard model for transitions, with a clock reset specification. The parameters of these exponential baseline hazards are jointly optimized with the model parameters.
Moreover, we impose a particular substructure on the model, where $\forall(k,k^{\prime})\in E,\,\beta^{k^{\prime}\mid k}=\beta$.
In the light of Figure 6, the regression and link functions were taken to be scaled and shifted hyperbolic tangent and its derivative respectively such that the normalized MMSE score always has a value of $1$ when $t\to$-$\infty$ and is not increasing
$\begin{cases}h(t,\psi)=\psi_{1}\tanh\left(\frac{\psi_{3}-t}{\psi_{2}}\right)+(1-\psi_{1}),\\ g(t,\psi)=\frac{\partial}{\partial t}h,\\ \psi=\begin{pmatrix}\sigma(\gamma_{1}+b_{1})\\ \exp(\gamma_{2}+b_{2})\\ \gamma_{3}+b_{3}\end{pmatrix},\quad\sigma(z)=\frac{1}{1+e^{-z}}.\end{cases}$
At last, an individual covariates correspond to a pair of binary variables indicating wheter or not the individual has a diploma and is a male $\left(\mathds{1}_{\text{CEP}}(i),\mathds{1}_{\text{male}}(i)\right)\in\{0,1\}^{2}$.
7.2 Results
7.2.1 Inference
Fitting was performed using $80\%$ of the data, with the remaining $20\%$ used for testing. The model was fitted using the Adam optimizer. The optimization process was run until convergence based on the same criterion as the simulation study with a tolerance of $5\%$, which was obtained after a little less than $500$ iterations with $10$ parallel chains.
Before displaying the fitted parameters, the Fisher Information Matrix was computed, implementing the method described in [baey2023efficient], directly taking into account shared parameters. The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix. The results are summarized in Table 8.
Parameter Inferred value Std. error Parameter Inferred value Std. error
$\gamma_{1}$ 2.1600 0.0470 $\gamma_{2}$ 0.1590 0.0140
$\gamma_{3}$ 1.0010 0.0030 $\tilde{Q}_{1}$ -0.5700 0.0430
$\tilde{Q}_{2}$ 0.9570 0.0960 $\tilde{Q}_{3}$ 0.3820 0.0430
$\tilde{Q}_{4}$ -0.2050 0.1570 $\tilde{Q}_{5}$ -2.3310 0.1380
$\tilde{Q}_{6}$ 0.5220 0.0470 $\tilde{R}$ 2.0830 0.0260
$\alpha^{3\mid 1}$ 0.7220 0.7150 $\alpha^{1\mid 0}$ -0.5400 0.4030
$\alpha^{2\mid 1}$ -1.0030 0.2000 $\alpha^{2\mid 0}$ -0.1210 0.0380
$\alpha^{3\mid 2}$ -0.0480 0.2520 $\alpha^{3\mid 0}$ -0.3070 0.0740
$\beta_{1}$ -0.2990 0.9130 $\beta_{2}$ 0.0970 1.0300
Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$ (variance parameters excluded).
Notably, the estimated standard errors are consistent with the number of observations per transition, with larger standard errors for transitions with fewer observations. Therefore, parameters with too few observations should be interpreted with caution, if at all. The linear covariate parameters also exhibit large standard errors, which indicates the coefficients are not significative in this setting.
The association parameters $\alpha$ being negative, we can infer that cognitive decline (i.e. rate of change) is associated with an increase in physical dependency.
7.2.2 Dynamic Prediction
Using the $n_{\text{test}}=100$ individuals not included in the model fitting, we performed dynamic prediction. For each individual, longitudinal measurements and trajectories were truncated at various times $t$, and predictions were made for future time points $u\geq t$ beyond the truncation. Accuracy was then assessed by comparing the most likely predicted state with the true state at each corresponding future time point $u$, as illustrated in Figure 9.
Formally, the prediction function $\chi_{u}^{*}$ is defined by Example 1, and the accuracy measure at time $t$ for predicted states $(s_{i})_{i}$ is defined as
$\text{accuracy}(u)=\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\mathds{1}_{\chi_{u\wedge C_{i}}(\mathcal{T}_{i})=s_{i}}.$
Although $\mathcal{T}_{i}^{*}$ is unkown, the quantity $\chi_{u\wedge C_{i}}(\mathcal{T}_{i})$ may be computed based on the censored trajectories. However, incorporating censoring times into prediction functions relies on information that is, in principle, not available at prediction time. Nevertheless, this inclusion is necessary to ensure fairness in accuracy metrics, since individuals who are not censored are generally more likely to have progressed to more advanced states of physical dependency than those who are censored.
Figure 9: Accuracy of the predicted state at each time point with respect to the true state for different truncation times.
Here, the initial drops correspond to the truncation times, before which the accuracy is always $100\%$.
As expected from dynamic predictions, the accuracy increases as more data is available, therefore when the truncation time grows. We also observe that the longer the prediction horizon, the lower the accuracy of the predictions.
8 Conclusion
We presented a general likelihood-based framework for joint modeling of nonlinear longitudinal biomarkers and multi-state survival processes. The model unifies both components on arbitrary directed graphs, extending classical joint models to accommodate Markov and semi-Markov structures, recurrent transitions, and nonlinear mixed-effects submodels. This formulation provides a flexible and rigorous foundation for linking longitudinal biomarker trajectories to complex event histories within a single coherent probabilistic model.
Simulation studies confirmed robust parameter recovery, and an application to the PAQUID cohort highlighted the ability to capture the interplay between cognitive decline and physical dependency.
Since the model may potentially involve many transitions and therefore a large number of parameters, a strategy of parameter sharing across transitions could help reduce the overall number of parameters. This is especially beneficial in settings with limited data for certain transitions, or when biological or clinical knowledge suggests similar effects across multiple transitions. Parameter sharing improves statistical efficiency, reduces overfitting, and facilitates model interpretability.
Future directions include a spline-based parametrization of baseline hazards to allow for nonparametric estimation. This approach is particularly useful when the true baseline hazard is expected to be complex or non-monotonic, or when parametric forms (such as exponential or Weibull) may be too restrictive and lead to model misspecification.
Another promising direction is the development of efficient visualization tools for multi-state trajectories and dynamic predictions. Interactive representations of individual event histories, transition probabilities, and longitudinal biomarker evolution would greatly facilitate model interpretation and communication of results to applied researchers.
Funding
This work was partially funded by the Stat4Plant project ANR-20-CE45-0012.
Acknowledgments
The authors would like to thank Jean-Benoist Leger for helpful discussions.
Conflict of interest
The authors declare that they have no conflicts of interest.
Data availability
The data analyzed in this study are publicly available via the R package timeROC (dataset Paquid).
Appendix A Multi-State Simulation
In nonlinear joint models with known design and known parameters parameters $\theta=(\gamma,Q,R,\alpha,\beta)$, the occurrence times of events conditionally on the latent variables $b_{i}$ and the covariates $X_{i}$ can be easily sampled using the inverse cumulative distribution function transform method. This may be achieved through bisection or other root-finding algorithms.
The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property by considering one transition at a time, according to Algorithm 2. The algorithme is similar to Gillespie’s algorithm [gillespie2007stochastic], and can also include a survival condition $T_{i1}\geq t_{i}^{\text{surv}}$ that uses the Chasles relation.
Algorithm 2 Simulation of trajectory $i$
1:subject-specific censoring time $C_{i}\in\mathbb{R}\cup\{+\infty\}$; covariates $X_{i}$; latent variables $b_{i}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i}^{\text{surv}}$, defaults to $t_{i}^{\text{surv}}=$-$\infty$.
2:
3:Initialize $\ell_{i}\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$.
4:
5:while $T_{i(\ell_{i}-1)}<C_{i}$ and $\{s:(S_{i(\ell_{i}-1)},s)\in E\}\neq\emptyset$ do
6:  for each $s$ with $(S_{i(\ell_{i}-1)},s)\in E$ do
7:   Draw $T_{i\ell_{i}}^{s}$ such that
8:      $\log\mathbb{P}(T_{i\ell_{i}}^{s}>t)=\int_{T_{i(\ell_{i}-1)}\vee t_{i}^{\text{surv}}}^{t}\lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}(w\mid b_{i},X_{i},T_{i(\ell_{i}-1)},\theta)\,dw,$
9:      $\forall t\geq T_{i(\ell_{i}-1)}\vee u.$
10:  end for
11:
12:  Set $T_{i\ell_{i}}\leftarrow\min_{m}T_{i\ell_{i}}^{s}$, and $S_{i\ell_{i}}\leftarrow\arg\min_{m}T_{i\ell_{i}}^{s}$.
13:  Append $(T_{i\ell_{i}},S_{i\ell_{i}})$ to $\mathcal{T}_{i}$.
14:  $\ell_{i}\leftarrow\ell_{i}+1$.
15:end while
16:
17:if $T_{i(\ell_{i}-1)}>C_{i}$ then
18:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$.
19:end if
20:return $\mathcal{T}_{i}$
The proof of the exactness of this algorithm is given below.
Proof.
First consider the case $t_{i}^{\text{surv}}=$-$\infty$. Let $t>T_{i\ell_{i}}$:
$\mathbb{P}(T_{i(\ell_{i}+1)}>t\mid T_{i\ell_{i}},S_{i\ell_{i}})$ $=\mathbb{P}(\bigcap_{s:\,(E_{i(\ell_{i}-1)},s)\in E}T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\prod_{s:\,(S_{i\ell_{i}},s)\in E}\mathbb{P}(T_{i(\ell_{i}+1)}^{s}>t\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
Furthermore:
$\mathbb{P}(S_{i(\ell_{i}+1)}\mid T_{i(\ell_{i}+1)}=t,T_{i\ell_{i}},S_{i\ell_{i}})$
$\propto p(\{T_{i(\ell_{i}+1)}^{S_{i(\ell_{i}+1)}}=t\}\cap\bigcap_{\begin{subarray}{c}s:\,(S_{i\ell_{i}},s)\in E,\\ s\neq S_{i(\ell_{i}+1)}\end{subarray}}\{T_{i(\ell_{i}+1)}^{s}\geq t\}\mid T_{i\ell_{i}},S_{i\ell_{i}}),$
$=\lambda_{i}^{S_{i(\ell_{i}+1)}\mid S_{i\ell_{i}}}\big(t\mid\theta,X_{i},T_{i\ell_{i}},b_{i})\exp\Big(-\sum_{s:\,(S_{i\ell_{i}},s)\in E}\Lambda_{i}^{s\mid S_{i\ell_{i}}}\big(t\mid b_{i},X_{i},T_{i\ell_{i}},\theta)\Big).$
Combining both, we retrieve the same joint density.
If $t_{i}^{\text{surv}}>$-$\infty$, one can also check that both densities are equal. Note that the condtioning only affects the first transition, as the next time $\forall\ell_{i}>0,\,T_{i\ell_{i}}\geq t_{i}^{\text{surv}}\implies\forall\ell_{i}>0,\,T_{i\ell_{i}}\vee t_{i}^{\text{surv}}=T_{i\ell_{i}}$.
Appendix B Expression of the Semi-Markov likelihood
Proof.
The proof is short and relies on the fact that, using the semi-Markov property 2
$p\left((T_{i(\ell+1)},S_{i(\ell+1)})_{\ell\leq m_{i}-1}\mid b_{i},X_{i},\theta\right)$ $=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell^{\prime}},S_{i\ell^{\prime}})_{\ell^{\prime}\leq\ell},\theta\right),$
$=\prod_{\ell=0}^{m_{i}-1}p\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell}),\theta\right).$
On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise
$\mathbb{P}(T_{i(m_{i}+1)}>C_{i}\mid\theta,X_{i},(T_{im_{i}},S_{im_{i}}),b_{i})=\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{s\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right),$
and so we recover the formula cited above with the convention that an empty sum is 0.
Appendix C Stochastic optimization process
(a) $\gamma$
(b) $Q$
(c) $R$
(d) $\alpha^{1\mid 0}$
(e) $\alpha^{2\mid 0}$
(f) $\alpha^{2\mid 1}$
(g) $\beta^{1\mid 0}$
(h) $\beta^{2\mid 0}$
(i) $\beta^{2\mid 1}$
Figure 10: Evolution of the parameters during the optimization of the marginal log-likelihood using stochastic gradient descent. The dotted lines correspond to the true values.