### v4

# A General Framework for Joint Multi-State Models

Abstract
Conventional joint modeling approaches generally characterize the relationship between longitudinal biomarkers and discrete event occurrences within terminal, recurring or competing risk settings, thereby offering a limited representation of complex, multi-state trajectories.
We propose a general multi-state joint modeling framework that unifies longitudinal biomarker dynamics with multi-state time-to-event processes defined on arbitrary directed graphs. The proposed framework also accommodates nonlinear longitudinal submodels and scalable inference via stochastic gradient ascent. This formulation encompasses both Markovian and semi-Markovian transition structures, allowing recurrent cycles and terminal absorptions to be naturally represented. The longitudinal and event processes are linked through shared latent structures within nonlinear mixed-effects models, extending classical joint modeling formulations.
We derive the complete likelihood, model selection criteria, and develop scalable inference procedures based on stochastic gradient ascent to enable high-dimensional and large-scale applications. In addition, we formulate a dynamic prediction framework that provides individualized state-transition probabilities and personalized risk assessments along complex event trajectories.
Through simulation and application to the PAQUID cohort, we demonstrate accurate parameter recovery and individualized prediction.
keywords:
joint modeling, multi-state processes, longitudinal data, survival analysis, stochastic gradient descent, dynamic prediction
1 Introduction
Joint modeling of longitudinal and time-to-event data has become an essential tool of modern biostatistics [papageorgiou2019overview], particularly for dynamic prediction in clinical applications. Classical joint models typically couple a longitudinal biomarker process with a single time-to-event outcome [rizopoulos2012joint], allowing for the integration of biological knowledge and individual heterogeneity via shared latent structures. However, many real-world processes involve multiple possible outcomes, intermediate stages, or recurrent events, which cannot be fully captured by a single-event framework [KrolMauguen2017]. To address these limitations, specialized frameworks have been developed to handle competing risks and recurrent events. However, these remain constrained to specific event types and fail to accommodate more general multi-state transitions. In such settings, multi-state models provide a more general and flexible approach [ferrer2016joint, you2024joint].
Multi-state models represent an overall joint process of clinical course as a series of discrete stages or health states that occur sequentially [lovblom2024modeling]. In biostatistics, these models are widely used for survival and reliability analysis, allowing for a richer and more accurate representation by capturing alternative paths to an event of interest, intermediate events, and progressive disease. Key components of multi-state models include transition intensity functions, which denote the instantaneous risk of moving from one state to another, and transition probability functions, which describe the probability of transition over longer intervals. Often, these models assume the Markov property, where future transitions depend only on the current state, simplifying the transition intensity functions [asanjarani2022estimation]. This assumption can be relaxed by adopting a semi-Markov formulation, in which the transition probabilities depend not only on the current state but also on the sojourn time, effectively resetting the time scale after each transition.
The link between multi-state models and joint models arises when a multi-state process is integrated as a component within a broader joint modeling framework. While multi-state models, such as those described in the “sequential state framework” primarily focus on movements between discrete states, joint models operate under a “parallel trajectory framework” that combines a longitudinal process with a time-to-event process.
The model proposed by Ferrer et al. [ferrer2016joint] exemplifies this link by presenting a joint model for a longitudinal process (e.g., Prostate-Specific Antigen (PSA) measurements) and a multi-state process (e.g., clinical progressions in prostate cancer). These two sub-models are interconnected by shared random effects, allowing the model to account for the correlation between a continuous longitudinal biomarker trajectory and acyclic discrete transitions between health states. Specifically, the previously proposed mechanism [ferrer2016joint, you2024joint] is based on binary censoring indicators for each state, implying that any given state can be entered at most once, therefore precluding the modeling of recurrent events.
The model we propose operates on an arbitrary directed graph, enabling representation of complex state transitions and recurrent events. We derive the complete likelihood for the nonlinear joint model, introduce an efficient stochastic approximation inference method, and develop dynamic prediction tools. To illustrate its practical relevance, we conduct both a simulation study and a synthetic biomedical case study.
This paper is organized as follows. Section 2 reviews the background on multi-state and joint modeling frameworks. Section 3.2 introduces the proposed general multi-state joint modeling framework, detailing the likelihood formulation and proposed model selection criteria. Section 4 describes the stochastic-gradient–based inference algorithm, while Section 5 presents the dynamic prediction framework. Section 6 reports simulation experiments that assess parameter recovery and convergence properties of the proposed model, whereas Section 7 applies the framework to data from the PAQUID cohort [Letenneur1994]. Finally, Section 8 concludes with perspectives and future research directions.
2 Background
Joint modeling of longitudinal and time-to-event data provides a unified statistical framework to analyze the interplay between continuous biomarker dynamics and event occurrence processes. In its classical form, this framework couples a mixed-effects model describing individual biomarker trajectories with a survival model for event times [rizopoulos2012joint]. Such models have become central in biomedical research, particularly for dynamic prediction and personalized risk assessment [papageorgiou2019overview]. However, these traditional joint models generally focus either on a single terminal event, competing, or recurrent events, limiting their ability to represent more complex event histories that include intermediate, recurrent, or competing events. To address these limitations, multi-state extensions have been developed, providing a natural framework for describing transitions between multiple clinical or biological states over time [ferrer2016joint, KrolMauguen2017].
2.1 Multi-State Markov Processes
A multi-state stochastic process is defined as a process $S(t)$ for $t\geq 0$, where $S(t)$ can take a finite number of values (states), often labeled $1,2,\dots,p$. Quantities of interest typically include the probability of being in a certain state at a given time and the distribution of first passage times.
A Markov process (or continuous-time Markov chain) is a specific class of multi-state models in which future transitions between states depend only upon the current state. This Markov property means the process is memoryless. A key consequence is that the duration spent in any state follows an exponential distribution, implying a constant hazard rate for leaving that state [jackson2011multi, putter2007competing].
However, the Markov assumption can be unrealistic in many real-world applications. For example, in the study of human sleep stages, sojourn times often do not follow an exponential distribution [dong2008non, roever2010statistical], and in the case of chronic diseases such as AIDS, the risk of disease progression can depend on the time elapsed since infection [andersen1999multi]. To address these limitations, semi-Markov processes (SMPs) were introduced, which allow for arbitrary distributions of sojourn times while retaining the Markov property for the embedded discrete-time chain. This flexibility makes SMPs suitable for modeling complex disease progression and patient recovery scenarios [putter2007competing].
2.2 Multi-State Semi-Markov Processes
Multi-state semi-Markov processes (MSMPs) offer a natural generalization by allowing the distribution of sojourn times to be arbitrary.
In an MSMP, the process is defined by a directed graph $G=(V,E)$, where $V$ represents the set of states and $E$ the set of possible transitions. The transition intensities $\lambda_{k\to k^{\prime}}(t\mid t_{0})$ from state $k$ to state $k^{\prime}$ at time $t$, given entry time $t_{0}$, are therefore assumed to be non identically zero. The sojourn time in state $k$ is not restricted to an exponential distribution, allowing for more realistic modeling of the time spent in each state.
MSMPs have been widely applied in a range of disciplines. In reliability, they are used to model degradation and repair processes [limnios2001semi]. In biomedical studies, they have proven useful for analyzing illness-death models or disease progression with non-exponential transitions [commenges2006likelihood, fiocco2008modelling]. Applications in finance also exist, where semi-Markovian dynamics can model credit rating migrations or economic regimes [luciano2006semi]. In these contexts, MSMPs retain the Markov property in the embedded jump chain while offering increased realism through flexible dwell time modeling.
From a methodological standpoint, MSMPs extend estimation strategies developed for Markov models, such as maximum likelihood or Bayesian inference, and can accommodate interval-censored or misclassified data [frydman2005estimation]. This makes them a powerful and general tool for multi-state event history analysis.
While both joint models and multi-state models offer powerful tools, they primarily address different facets of disease progression and have complementary strengths. Multi-state models excel at understanding the sequence and timing of discrete events, while joint models are adept at modeling continuous biomarker trajectories and their associations with event outcomes for dynamic prediction. The complexity of real-world diseases often necessitates a more integrated approach that can leverage the strengths of both frameworks. For instance, in prostate cancer, multiple types of relapse may occur successively (a multi-state process), and their risk is influenced by the dynamics of longitudinal biomarkers like PSA.
2.3 Joint Models and Multi-State Processes
Previous research has initiated such integration, with models like the one proposed by Ferrer et al. [ferrer2016joint], which combines a linear mixed sub-model for longitudinal data with a multi-state sub-model using shared random effects. This model enables the simultaneous analysis of repeated measurements of a biomarker (the longitudinal process, e.g., PSA levels) and the times of transitions between multiple health states (the multi-state process). It explicitly accounts for the link between these two correlated processes and uses information from the biomarker dynamics to explain or predict clinical progression events. The multi-state sub-model assumes a non-homogeneous Markov process for the clinical progression, meaning that the future of the process depends only on the current state (Markov property) and the time elapsed since study entry (non-homogeneous property).
While both joint models and multi-state models offer powerful tools for analyzing disease progression, they emphasize different aspects of the underlying process. Multi-state models capture the sequence and timing of discrete transitions between states, whereas joint models focus on the evolution of continuous longitudinal markers and their association with event risk. Existing joint multi-state models, such as that of Ferrer et al. [ferrer2016joint], are restricted to linear mixed-effects submodels and to acyclic Markov transition graphs, since their reliance on binary censoring indicators inherently prevents modeling recurrent events. However, many biological processes involve cyclic or semi-Markov transitions and nonlinear biomarker dynamics. These cannot be handled by existing frameworks, motivating the present general formulation.
In contrast, many real-world processes—such as recurrent disease relapse or recovery—require more flexible representations involving nonlinear biomarker dynamics, semi-Markovian dependencies, and recurrent cycles. The next section introduces a general framework that integrates these features within a unified likelihood-based formulation, enabling efficient inference and dynamic prediction along complex event trajectories.
3 A General Multi-State Joint Modeling Framework
To address the limitations imposed by traditional joint modeling frameworks, we extend them to incorporate a multi-state process, thereby enabling a unified analysis of longitudinal biomarkers and discrete transitions between multiple health states.
The proposed model defines a joint process $\{Y_{i}(t),S_{i}(t)\}_{t}$, where $Y_{i}(t)$ denotes the longitudinal biomarker trajectory and $S_{i}(t)$ the multi-state event process. Both components are linked through shared latent structures, and the event process is represented on an arbitrary directed graph supporting both Markovian and semi-Markovian transition mechanisms.
3.1 Notation Conventions
Throughout this section, let $i\in\llbracket n\rrbracket$ index $n\geq 1$ individuals, and let $t_{ij}$ denote the $j$-th observation time of individual $i$. The vector $Y_{ij}\in\mathbb{R}^{d}$ represents longitudinal biomarker measurements, and $X_{i}\in\mathbb{R}^{p}$ the associated covariates.
The latent random effects $b_{i}\sim\mathcal{N}(0,Q)$ capture individual heterogeneity, $\gamma$ represent shared characteristics across all individuals, and the subject-specific parameters are defined as $\psi_{i}\coloneqq f(\gamma,X_{i},b_{i})$. The set $\mathcal{H}_{i}(t)\coloneqq\left\{(\sigma,h(s,\psi_{i})):s\leq t\right\}$ represents the true underlying history of the longitudinal process up to time $t$. Correspondingly, $\mathcal{Y}_{i}(t)$ denotes the observed noisy longitudinal history, and $\mathcal{S}_{i}(t)$ denotes the observed trajectory of the semi-Markovian dynamics.
Transition times and states are denoted by $(T_{i\ell},S_{i\ell})\in\mathbb{R}^{+}\times V$, where $V$ is the space of admissible states, and are subject to right-censoring at time $C_{i}\in\bar{\mathbb{R}}^{+}$. We define the index of the last observed transition before time $t$ as $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$. The complete sequence of transition times and states for individual $i$ is denoted by $\mathcal{T}_{i}^{*}$, whereas $\mathcal{T}_{i}$ denotes the corresponding right-censored sequence.
3.2 Multi-State Processes on Arbitrary Graphs
The joint model proposed by Ferrer et al. [ferrer2016joint] established a foundational link between longitudinal biomarkers and multi-state processes through shared random effects. However, their formalism relies on binary censoring indicators to characterize transitions, thereby inherently restricting the modeling framework to Directed Acyclic Graphs (DAGs) and excluding recurrent events. Indeed, the proposed model uses, for each state $k\in E$, a tuple $(T_{k},\delta_{k})$, where $T_{k}$ denotes the observed or censored time until leaving state $k$ and $\delta_{k}$ is an indicator equal to $1$ if the transition was observed and 0 if the observation was censored (i.e., the transition was not observed).
To overcome this limitation, we present a new method based on arbitrary graph transitions. By shifting away from binary indicators toward a more flexible graph-theoretic specification, our approach accommodates complex disease trajectories, including recurrent events and cyclic transitions that cannot be captured within a DAG-based structure.
3.2.1 Graph Structure
Let $G=(V,E)$ be a directed graph, where $V=\llbracket p\rrbracket$ denotes the set of states and $E\subseteq V\times V$ the set of admissible transitions. The graph encodes all possible paths of the multi-state process, allowing for competing, recurrent, or absorbing transitions. Figure 1 illustrates an example of a four-state model, including an absorbing state (Death). The corresponding adjacency matrix $A=(A_{kk^{\prime}})_{1\leq k,k^{\prime}\leq p}$ satisfies $A_{kk^{\prime}}=1$ if $(k,k^{\prime})\in E$ and 0 otherwise.
Figure 1: Example of a 4-state transition graph $G=(V,E)$, including an absorbing state (4).
Such a representation generalizes standard joint models, which can be recovered as special cases: single-event (linear chain), competing-risks (two absorbing transitions), or recurrent-event settings (cyclic transitions) as illustrated in Figure 2.
Standard Risk
Competing Risks
Recurrent Risk
Figure 2: Illustration of classical joint modeling approaches: (left) standard risk, (middle) competing risks, and (right) recurrent risk.
Such joint modeling offers a flexible framework that extends beyond classical approaches and is particularly useful in medical applications.
3.2.2 Individual Trajectories
Each individual $i\in\llbracket n\rrbracket$ follows a latent trajectory
$\mathcal{T}_{i}^{*}\coloneqq\left((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\right),$
where $T_{i\ell}\in\mathbb{R}^{+}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state. The observed trajectory is right-censored at time $C_{i}\in\bar{\mathbb{R}}^{+}$, so that only $\mathcal{T}_{i}\coloneqq\left\{(T_{i\ell},S_{i\ell}):T_{i\ell}\leq C_{i}\right\}$ is observed. Between transitions, the state $S_{i}(t)$ is constant, thus the transitions are the discontinuity points of the map $t\mapsto\mathcal{S}_{i}(t)$, of which we assume there are at most only infinitely countably many.
Letting $m_{i}(t)\coloneqq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$ be the index of the last observed transition before time $t$ and $m_{i}\coloneqq m_{i}(C_{i})$, we thus assume that no further transitions occur for individual $i$ on the interval $(T_{im_{i}},C_{i}]$. We also define $\mathcal{S}_{i}(t)\coloneqq\left\{(s,S_{i}(s)):s\leq t\right\}$ as the observed trajectory up to time $t$, which is equivalent to specifying all transition times and their corresponding states up to $t$, together with the time $t$ itself.
We emphasize once more that modeling the observed data solely as censored collections of transition times, as in Ferrer et al. [ferrer2016joint], is not sufficient to represent such trajectories, since it implicitly restricts each state to be visited at most once and therefore rules out recurrent transitions.
The process is assumed to satisfy a (possibly time-inhomogeneous) semi-Markov property
$\mathscr{L}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid\{(T_{i\ell^{\prime}},S_{i\ell^{\prime}})\}_{\ell^{\prime}\leq\ell}\right)=\mathscr{L}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid(T_{i\ell},S_{i\ell})\right),$ (1)
which is relaxed compared to a stricter Markov property by allowing sojourn-time dependence.
3.2.3 Transition Intensities and Survival Functions
Although one could equivalently define intensities for all ordered pairs of states and set those corresponding to inadmissible transitions identically to zero, explicitly restricting attention to a graph of admissible transitions provides a clearer and more coherent representation of the underlying state dynamics.
For any admissible transition $(k,k^{\prime})\in E$ with nonzero probability, let $\lambda_{i}^{k\to k^{\prime}}(t_{0},t)$ denote the instantaneous risk at time $t$, given entry time $t_{0}\leq t$, and conditionally the transition $k\to k^{\prime}$
$\lambda_{i}^{k\to k^{\prime}}(t_{0},t)\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(\ell+1)}\leq t+\delta\mid T_{i(\ell+1)}>t,T_{i\ell}=t_{0},S_{i\ell}=k,S_{i(\ell+1)}=k^{\prime})}{\delta}.$
Assuming this quantity to be finite for any pair $(t_{0},t)$ so that the underlying distribution of transition times is continuous, the cumulative risk and the corresponding survival function are
$\Lambda_{i}^{k\to k^{\prime}}(t_{0},t)=\int_{t_{0}}^{t}\lambda_{i}^{k\to k^{\prime}}(t_{0},s)\,ds,$
$\forall t\geq T_{i\ell},\,\mathbb{P}(T_{i(\ell+1)}>t\mid T_{i\ell},S_{i\ell})=\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},t)\right).$
Conditional transition probabilities are
$\mathbb{P}(S_{i(\ell+1)}=k^{\prime}\mid T_{i(\ell+1)},T_{i\ell},S_{i\ell})=\frac{\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},T_{i(\ell+1)})}{\sum_{j:(S_{i\ell},j)\in E}\lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})},$
so the joint density of a transition writes as
$p\left((T_{i(\ell+1)},T_{i(\ell+1)})\mid(T_{i\ell},T_{i\ell})\right)=$ $\frac{\lambda_{i}^{S_{i\ell}\to S_{i(\ell+1)}}(T_{i\ell},T_{i(\ell+1)})}{\sum_{j:(S_{i\ell},j)\in E}\lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})}$
$\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},T_{i(\ell+1)})\right)$ (2)
In practice, numerical integration can be efficiently carried out using quadrature methods such as the Gauss-Legendre quadrature [lether1978construction].
3.3 Multi-State Joint Model Specification
Omitting the implicit conditioning on the individual parameters and covariates for the sake of simple notation, the proposed multi-state joint model with a Gaussian prior and homoscedastic Gaussian noise is specified by
$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \lambda_{i}^{k\to k^{\prime}}(t_{0},t)=\lambda_{0}^{k\to k^{\prime}}(t_{0},t^{\prime})\exp\left(\alpha^{k\to k^{\prime}}g^{k\to k^{\prime}}(t,X_{i},\psi_{i})+\beta^{k\to k^{\prime}}X_{i}\right),\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q).\end{cases}$
Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k\to k^{\prime}}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k\to k^{\prime}}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$. This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications.
The coefficient $\alpha^{k\to k^{\prime}}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$. More concretely, for a one-unit increase in $g^{k\to k^{\prime}}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k\to k^{\prime}})$. Likewise, $\beta^{k\to k^{\prime}}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k\to k^{\prime}}_{j})$. Thus, $\alpha^{k\to k^{\prime}}$ captures the degree to which the biomarker trajectory (via its latent parameters $\psi_{i}$ or function $g^{k\to k^{\prime}}$) influences the transition risk, while $\beta^{k\to k^{\prime}}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.
3.3.1 Model Variants
Two baseline hazard conventions are common and can be specified as
$\lambda_{0}^{k\to k^{\prime}}(t_{0},t)=\begin{cases}\lambda_{0}^{k\to k^{\prime}}(t_{0},t)&\text{(clock-reset)},\\ \lambda_{0}^{k\to k^{\prime}}(t)&\text{(clock-forward)}.\end{cases}$
The clock-reset form models risk as a function of time spent in the current state, whereas the clock-forward form measures risk with respect to global time since study entry. Common parametric baseline hazard specifications include the exponential ($\lambda_{0}(t)=\lambda$), the Weibull ($\lambda_{0}(t)=k\lambda^{k}t^{k-1}$), and the Gompertz ($\lambda_{0}(t)=ae^{bt}$) families, among others.
Furthermore, under the sojourn-time formulation, transition times are no longer constrained to be positive, since only the elapsed duration between successive transitions enters the baseline hazard function.
3.3.2 Likelihood Formulation
To derive the marginal likelihood, we decompose the joint distribution of the longitudinal and multi-state processes under a set of standard conditional independence assumptions detailed in Table 3.3.2 below.
Assumptions for Likelihood Factorization A. Latent-level independence A1. Random effects $(b_{i})_{i}$ are mutually independent across individuals. B. Conditional independence within individuals B1. Longitudinal observations $(Y_{ij})_{ij}$ are mutually independent given $b_{i},X_{i}$. B2. Trajectories $(\mathcal{T}_{i}^{*})_{i}$ are mutually independent given $b_{i},X_{i}$. B3. Conditional on $b_{i},X_{i}$, the longitudinal and event processes are mutually independent. C. Censoring and process assumptions C1. Censoring times $(C_{i})_{i}$ are mutually independent and noninformative, i.e., $C_{i}\mathrel{\perp\!\!\!\perp}(Y_{i},\mathcal{T}_{i}^{*})\mid b_{i},X_{i}$. C2. Conditionally on $b_{i},X_{i}$, event trajectories satisfy the semi-Markov property 1 given $b_{i},X_{i}$.
These grouped assumptions mirror those used in classical joint modeling [rizopoulos2012joint] and multi-state survival analysis [putter2007competing]: independence across subjects, the conditional independence structure linking the longitudinal and event submodels through shared random effects, noninformative censoring, and semi-Markovian dynamics.
Let $\theta\coloneqq(\gamma,Q,R,\alpha,\beta)$ denote the vector of model parameters. For subject $i\in\llbracket n\rrbracket$, we observe the longitudinal measurements $Y_{i}=(Y_{i1},\dots,Y_{in_{i}})$ and the event trajectory $\mathcal{T}_{i}=\left\{(T_{i\ell},S_{i\ell})_{\ell\leq m_{i}}\right\}$. The joint likelihood then factorizes as
$p_{\theta}(Y_{i},\mathcal{T}_{i}\mid X_{i},C_{i})=\int p_{\theta}(Y_{i}\mid b_{i},X_{i})\,p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i})\,p_{\theta}(b_{i})\,db_{i}.$
This formulation generalizes the joint likelihoods of wulfsohn1997joint and ferrer2016joint to arbitrary multi-state event structures. Each part can then be explicitly expressed using Assumptions 3.3.2, yielding an expression very similar to that obtained by rizopoulos2012joint.
Prior likelihood:
$p_{\theta}(b_{i})=\frac{(2\pi)^{-q/2}}{\det(Q)^{1/2}}\exp\left(-\tfrac{1}{2}b_{i}^{T}Q^{-1}b_{i}\right),$
Longitudinal likelihood:
$p_{\theta}(Y_{i}\mid b_{i},X_{i})=\prod_{j=1}^{n_{i}}\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
$\text{with }\psi_{i}=f(\gamma,X_{i},b_{i}).$
Semi-Markov likelihood:
$p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)=$ $\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right)$ (3)
$\exp\left(-\sum_{j:(S_{im_{i}},j)\in E}\Lambda_{i,\theta}^{S_{im_{i}}\to j}(T_{im_{i}},C_{i}\mid b_{i},X_{i})\right).$
The proof of the expression of the Semi-Markov likelihood (3) is provided in Appendix B. Each term $p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right)$ can be computed using Equation 2.
The likelihood factorization above is valid and forms the basis for scalable inference procedures described in Section 4, leveraging the complete trajectory of biomarkers to refine predictions of transitions and survival [de2010mstate].
Moreover, here, we assume that the initial state $S_{i0}$ is observed. However, the framework could also incorporate a multinomial model for unobserved initial states [yiu2018clustered].
4 Statistical Inference and Model Selection
In this section, we outline a practical framework for statistical inference in the proposed multi-state semi-Markov joint model. We first derive an optimization scheme for estimating model parameters through a stochastic gradient ascent procedure, and then introduce model selection criteria to assess competing specifications. These methods will be illustrated and evaluated on simulated and real datasets in Section 6.
4.1 Stochastic Gradient Ascent
The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\mathrm{complete}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\mathrm{marginal}}$, obtained by integrating them out
$\mathcal{L}_{\mathrm{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\mathrm{complete}}(\theta;X,Y,\mathcal{T},C,b)\,db,$
where $b\coloneqq(b_{1},\dots,b_{n})$, and owing to the independence across individuals
$\mathcal{L}_{\mathrm{complete}}(\theta;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p_{\theta}(Y_{i}\mid b_{i},X_{i})\,p_{\theta}(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i})\,p_{\theta}(b_{i}).$
It is to be noted that this integral is typically intractable in closed form due to the nonlinear nature of the model. Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace] and Gauss–Hermite quadrature, alongside MCMC-based approximations such as Metropolis–Hastings [hastings1970monte], Metropolis-within-Gibbs, Hamiltonian Monte Carlo [neal2011mcmc]. These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic].
Under interchangeability of integral and differentiation, setting $x\coloneqq(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
$\nabla_{\theta}\log\mathcal{L}_{\mathrm{marginal}}(\theta;x)=\mathbb{E}_{b\sim p_{\theta}(\cdot\mid x)}\left[\nabla_{\theta}\log\mathcal{L}_{\mathrm{complete}}(\theta;x,b)\right].$
This expectation is approximated using Monte Carlo samples from the posterior $p_{\theta}(\cdot\mid x)$, avoiding the need to evaluate the intractable marginal likelihood $\mathcal{L}_{\mathrm{marginal}}(\theta;x)$. To ensure the covariance matrices $Q$ and $R$ remain symmetric and positive-definite during optimization, we parametrize them using a $\log$-Cholesky decomposition of their inverse. By taking the logarithm of the diagonal elements of the Cholesky factors, we map the constrained covariance parameters to an unconstrained real-valued space, and avoid explicit matrix inversion for likelihood computation. This formulation naturally supports stochastic gradient ascent with Markov Chain Monte Carlo (MCMC)-based posterior sampling and enables parallelization across individuals, ensuring convergence to a critical point. The update rule is detailed in Algorithm 1.
Algorithm 1 Stochastic gradient ascent model inference
1:data $x$; initial parameters $\theta^{(0)}$; step sizes $(\eta_{t})_{t\geq 0}$ with $\sum_{t\geq 0}\eta_{t}=+\infty$, $\sum_{t\geq 0}\eta_{t}^{2}<+\infty$; number of parallel chains $K$; MCMC samplers $\{p_{k}^{\mathrm{MCMC}}\}_{k=1}^{K}$; stopping criterion.
2:
[…]
4:while not converged do
5:  For each $k\in\llbracket K\rrbracket$, draw $b_{k}\sim p_{\theta^{(t)}}(\cdot\mid x)\approx p_{k}^{\mathrm{MCMC}}$.
6:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\frac{\eta_{t}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}_{\mathrm{complete}}(\theta^{(t)};x,b_{k})$
7:  $t\leftarrow t+1$
8:end while
9:return $\hat{\theta}$
Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient ascent framework, or even specialized optimizers such as Adam [kingma2014adam], NAdam [dozat2016incorporating] or Adagrad [duchi2011adaptive]. In particular, the Adam optimizer will later be used in our numerical experiments.
4.2 Model Selection and Information Criteria
To compare competing model specifications, one typically relies on information criteria based on the marginal likelihood. In practice, competing models are compared by selecting the specification that yields the smallest value of the chosen information criterion. For instance, the Akaike Information Criterion (AIC) [akaike2003new] and the Bayesian Information Criterion (BIC) [schwarz1978estimating] are standard tools. While the AIC relies on an asymptotically unbiased estimator of the $\log$-likelihood, the BIC aims to identify the true model with high probability as the sample size grows.
The Akaike Information Criterion is computed as follows
$\mathrm{AIC}\coloneqq-2\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2\dim\hat{\theta}$,
where $\dim\hat{\theta}$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
The Bayesian Information Criterion is similarly defined but imposes a stronger penalty for model complexity, growing in $\log n$ where $n$ is the number of observations. Particularly in the context of mixed effects models, and even more so for multi-state joint models, the number of observations $n$ may vary from one definition to another, either the total number of repeated measurements or the number of individuals [delattre2014note]. To alleviate this problem, we recall the derivation of the BIC from the Laplace approximation [lebarbier2006introduction]. For a prior distribution $\pi$ on a set of models $\mathcal{M}$, the Laplace approximation yields, up to constant terms (which could also be included to improve the approximation), the approximation of the posterior probability given observed data $x$ reads
[…]
where $H_{\hat{\theta}}$ denotes the Hessian matrix of the marginal $\log$-likelihood evaluated at the Maximum Likelihood Estimator (MLE). Then, under standard regularity conditions that allow differentiation and integration to be interchanged twice, we have that
[…]
Thus, our BIC criterion can be written as
[…]
However, as previously noted, the marginal $\log$-likelihood
$\log\mathcal{L}_{\mathrm{marginal}}(\theta;x)=\log\int\mathcal{L}_{\mathrm{complete}}(\theta;x,b)\,db$
is generally intractable in mixed effects and joint models, since the integral with respect to the latent variables does not admit a closed form. Although the gradient of the marginal $\log$-likelihood can be efficiently estimated using Fisher’s identity, the marginal $\log$-likelihood itself cannot be evaluated directly.
One possibility is to estimate it using bridge sampling, which has been proven to provide consistent and accurate estimators of normalizing constants and marginal likelihoods based on MCMC output [meng1996simulating, gronau2017tutorial].
An alternative consists in approximating the posterior distribution $p_{\hat{\theta}}(b\mid x)$ by a Gaussian distribution, for instance via a Laplace approximation. Using the entropy identity,
$\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};x)=\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log\mathcal{L}(\hat{\theta};x,b)\right]-\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log p_{\hat{\theta}}(b\mid x)\right],$
we can further leverage independence across individuals, so that the entropy term decomposes as
$\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log p_{\hat{\theta}}(b\mid x)\right]=\sum_{i=1}^{n}\mathbb{E}_{b_{i}\sim p_{\hat{\theta}}(\cdot\mid x_{i})}\left[\log p_{\hat{\theta}}(b_{i}\mid x_{i})\right].$
Approximating each posterior $p_{\hat{\theta}}(\cdot\mid x_{i})$ by a Gaussian distribution with covariance matrix $\Sigma_{i}$, we obtain
$\log\mathcal{L}_{\mathrm{marginal}}(\hat{\theta};x)\approx\mathbb{E}_{b\sim p_{\hat{\theta}}(\cdot\mid x)}\left[\log\mathcal{L}(\hat{\theta};x,b)\right]+\sum_{i=1}^{n}\underbrace{\frac{1}{2}\left(\log\det\Sigma_{i}+q\left(\log(2\pi)+1\right)\right)}_{H_{i}},$
where $q$ denotes the dimension of the latent vector.
This approximation replaces the intractable posterior entropy with a sum of closed-form entropies of $q$-dimensional Gaussian distributions. Furthermore, each term $H_{i}$ depends only on its corresponding covariance matrix $\Sigma_{i}$, which can be easily estimated from MCMC samples of the latent variables using the empirical covariance estimator. The practical relevance of the proposed model selection criteria is discussed in Subsection 6.3.
5 Dynamic prediction
In this section, we turn to dynamic prediction within the multi-state framework. Although the prediction of longitudinal trajectories does not fundamentally differ from the standard nonlinear joint modeling setting, the key difference arises from the structure of the event process. Indeed, because the joint distribution of multi-state trajectories has no tractable closed-form expression, direct calculation of predictive quantities is generally infeasible.
To address this, we rely on simulation-based procedures which are rendered possible using Algorithm 3 (detailed in the Appendix). Specifically, predictions are derived from simulated event trajectories, allowing for individualized risk assessment and forecasting of future states or event times based on a subject’s observed biomarker and event history. This approach generalizes classical dynamic prediction by leveraging the rich structure of multi-state processes and the joint distribution of longitudinal and event data.
Let $i$ index a new individual. Suppose we are interested in some functional of the true (unobserved) future trajectory,
$\chi^{*}(\mathcal{T}_{i}^{*})\in\mathbb{Y},$
where $\mathbb{Y}$ is some generic vector space.
For example, one can think of the state taken by the individual $i$ at a time $u$. As is the case in traditional dynamic prediction however, we are only interested in predicting quantities or characteristics that depend on a yet unobserved future, given prior information, i.e., the trajectory $\mathcal{S}_{i}(t)$ up to some prediction time $t\leq u$ and observed longitudinal markers $\mathcal{Y}_{i}(t)\coloneq\left\{(t_{ij},Y_{ij}):t_{ij}\leq t\right\}$. Contrary to (single transition) joint models, the conditional probability distribution on $(\mathbb{R}^{+}\times V)^{\mathbb{N}}$ cannot be analytically derived.
Nonetheless, as seen in Algorithm 3, we are able to accurately simulate each trajectory. Therefore, a single or double Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its conditional expected value (or other statistics depending on the posterior distribution, such as the median)
$\hat{\chi}_{i}^{(1)}\coloneqq\mathbb{E}_{\hat{\theta}}\left[\chi^{*}(\mathcal{T}^{\prime}_{i})\mid\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right]\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\mathrm{i.i.d.}}{\sim}\mathscr{L}_{\hat{\theta}}\left(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{S}_{i}(t)\right),$
where the samples $\{\mathcal{T}_{i}^{(k)}\}_{k=1}^{B}$ are generated using a MCMC algorithm targeting the distribution $\mathscr{L}_{\hat{\theta}}\left(b_{i}\mid\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right)$ in conjunction with Algorithm 3.
Another approach from Bayesian perspective treats the model parameters $\theta$ as random variables by imposing a prior distribution together with the training data $\mathcal{D}$ [rizopoulos2011dynamic], and thus in this case our new estimator may be defined
$\hat{\chi}_{i}^{(2)}\coloneqq\mathbb{E}_{\theta^{\prime}\sim\mathscr{L}(\theta\mid\mathcal{D})}\left[\mathbb{E}_{\theta^{\prime}}\left[\chi^{*}(\mathcal{T}^{\prime}_{i})\mid\mathcal{Y}(t),\mathcal{S}_{i}(t)\right]\right],$
and approximated by a double Monte Carlo scheme accordingly where $\mathscr{L}(\theta\mid\mathcal{D})$ is the posterior distribution of $\theta$.
In the case of longitudinal values, the procedure simplifies: for each draw of $b_{i}$, one usually predicts a single deterministic functional by substituting $\chi$ with $h(\cdot,\psi_{i})$, which already corresponds to the conditional mean of $Y_{i}$ given $b_{i}$. However, in general multi-state prediction, $\chi^{*}$ may very well depend on (countably) infinitely many transitions. As a result, since the proposed estimation relies on simulated samples, we require the quantity to depend only on a finite subset of transitions.
Assumption 1.
There exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}^{+}\times V)^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}\coloneqq\sigma\left((T_{il},S_{il})_{\ell\leq n}\right)$ such that
[…]
Essentially, given Assumption 1, with probability one we are guaranteed that the quantity of interest may be computed in a finite number of simulation steps. Indeed, the prediction algorithm for a single sample may be summarized as in Algorithm 2.
Algorithm 2 Prediction algorithm
[…]
2:
[…]
4:while $\ell\leq\tau_{i}$ do
5:  Append simulated $(T_{i\ell},S_{i\ell})$ to $\mathcal{T}_{i}$
6:  $\ell\leftarrow\ell+1$
7:end while
8:return $\chi\left((T_{il},S_{il})_{\ell\leq\tau_{i}}\right)$
Numerous quantities of interest may be encompassed by this framework. We give multiple examples below. First, for a directed acyclic graph $G=(V,E)$, we define the depth of $G$, denoted by $\operatorname{depth}(G)$, as the length of the longest directed path in $G$, that is,
$\operatorname{depth}(G)\coloneqq\max_{(v_{0},\dots,v_{k})}k,$
where $(v_{0},\dots,v_{k})$ ranges over all directed paths in $G$.
Example 1 (State at time $u$).
Let $G=(V,E)$ be a finite directed acyclic graph, $u\in\mathbb{R}^{+}$ be a fixed time, and $\mathbb{Y}=\Delta(V)$ be the simplex of distributions over states. Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq u\text{ or }S_{in}\text{ is absorbing}\}$. Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$. For each state $k\in V$, let $\mathds{1}_{k}\in\mathbb{Y}$ denote the vector that is $1$ at entry $k$ and 0 elsewhere (a one-hot encoding of $k$). Then, we set $\chi_{u}^{*}(\mathcal{T}_{i}^{*})=\mathds{1}_{S_{i\tau_{i}}}\in\mathbb{Y}$ to represent the state occupied by individual $i$ at time $u$. In particular, if $V=\{0,1\}$ and $E=\left\{(0,1)\right\}$, we recover the special case used for survival probability estimation in standard joint models [rizopoulos2011dynamic].
Example 2 (Hitting time).
Let $G=(V,E)$ be a finite directed acyclic graph, $\mathbb{A}\subseteq V$ a non-empty subset of states, and $\mathbb{Y}=\bar{\mathbb{R}}^{+}$. For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subseteq V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$. Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\vee\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$. Then, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$ and we set $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})=T_{i\tau_{i}}+\mathds{1}_{\{S_{i\tau_{i}}\}\not\rightsquigarrow\mathbb{A}}(+\infty)$ represents the hitting time for the set $\mathbb{A}$.
In Example 1, the stopping time $\tau_{i}$ captures the step at which the process for individual $i$ reaches or exceeds a given time $u$, or enters an absorbing state. The resulting value $\chi^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$ therefore represents the state of the individual at that specific time.
In Example 2, the stopping time $\tau_{i}$ corresponds to the first time the process reaches a target subset of states $\mathbb{A}$, or becomes unable to reach it in the future. The associated value $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})$ thus represents the hitting time of the set $\mathbb{A}$, which may be finite if $\mathbb{A}$ is reached, or $\infty$ otherwise.
Other practical applications may include finding the expected number of edges in some given trajectory, the number of times an individual has returned in a specific state, the expected time between transitions…
This simulation-based dynamic prediction framework generalizes classical approaches from single-event joint models to the multi-state context. It enables individualized forecasting of future state occupancy, event risks, sojourn times, and other clinically relevant outcomes, fully exploiting the subject’s observed biomarker trajectory and event history. In Section 7, we demonstrate its application to forecasting dependency trajectories in the PAQUID cohort (see Subsection 7.3.2).
6 Simulation Study
The purpose of this simulation study is to evaluate the finite-sample performance and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known. We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model. This setup allows us to assess both statistical estimation error (see Table 5) and computational scalability of the stochastic gradient estimation procedure. We further evaluate the model selection performance of the proposed methodology, as outlined in Subsection 4.2. To ensure statistical reliability, we perform $n_{\mathrm{runs}}=100$ independent simulations on datasets, comprised of $n=500$ individuals, generated from the same underlying distribution.
All simulations, estimations, and figures reported in this section were produced using the open-source jmstate Python package version 0.15.2, available on PyPI, which we developed to implement the multi-state joint modeling framework and inference algorithms described in Sections 3.2 and 4.
6.1 Model Specification
The simulated model represents a toy example based on a pharmacological scenario where drug absorption can influence seizure activity and patient stability. The model involves two states: Seizure ($1$) and Stable ($2$), which is absorbing. Transitions occur both from $1\rightarrow 1$ and $1\rightarrow 2$ (Figure 3), reflecting how a patient may either experience repeated seizure episodes or stabilize after treatment. If the drug is well absorbed, it helps mitigate seizure activity and promotes stabilization. Conversely, poor absorption increases the likelihood of recurrent seizures, highlighting the importance of pharmacokinetics in disease control. Transition intensities follow an exponential baseline hazard.
Figure 3: State transition diagram of the simulated recurrent two-state model.
The longitudinal biomarker trajectory is modeled using a bi-exponential function:
$h(t,\psi)=\psi_{1}\left(e^{-\psi_{2}t}-e^{-\psi_{3}t}\right),\quad\psi_{i}=\gamma e^{b_{i}},\quad b_{i}\sim\mathcal{N}\left(0,\mathrm{diag}(Q_{1},Q_{2},Q_{3})\right),$
and the observation noise is assumed Gaussian, independent across individuals and time points. This bi-exponential structure is commonly used in pharmacokinetic and longitudinal modeling (e.g., kerioui2022modelling), although alternative parametrizations are sometimes employed. Incidentally, one of the central points in kerioui2022modelling was the deliberate exclusion of recurrent transitions, regarded as difficult to estimate reliably. We show empirically that such transitions can in fact be estimated within our framework.
The link functions are applied consistently across all transitions and defined as
$g^{1\to 1}(t,\psi)=g^{1\to 2}(t,\psi)=\int_{0}^{t}h(s,\psi)\,ds,$
integrating the current biomarker value, thus providing a cumulative measure. Covariates $X_{i}$ are normally distributed and enter linearly via $\beta^{k\to k^{\prime}}X_{i}$.
Longitudinal data were collected at $20$ equally spaced time points between $t=0$ and $t=15$ per subject, with censoring at $C_{i}\sim\mathcal{U}([10,15])$ alongside the individual trajectories.
Since the base hazard rates $\lambda^{1\to 1}$ and $\lambda^{1\to 2}$ are jointly optimized with the other parameters, the final parameter vector lies in $\mathbb{R}^{13}$, and its values used in simulation are shown in Table 5. Each simulated dataset comprised a moderate number of $n=500$ individuals, generating roughly balanced transition counts (Figure 4). Thanks to careful parametrization, almost all parameters are mapped to the whole real axis unconstrained.
A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below. Note that unlike traditional joint modeling, the notion of censored transitions does not apply here. If a transition does not occur within a trajectory $\mathcal{T}_{i}$, it is simply omitted from the representation.
Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions histograms for the complete population. $76$ patients experienced repeated seizure.
6.2 Estimator Convergence
For each run, model parameters were estimated using the stochastic gradient ascent procedure described in Section 4, using a Metropolis-within-Gibbs sampling algorithm. Although the initial parameter values can strongly influence the optimization process, we set $\gamma^{(0)}=(1,1,1)^{T}$ to prevent division by zero, initialize both covariance matrices $Q$ and $R$ as identity matrices, set $\lambda^{1\to 1}$ and $\lambda^{1\to 2}$ to $1$, and initialize all linear coefficients to zero, thereby avoiding any assumptions about prior knowledge of the parameters.
We used the Adam optimizer alongside an adaptive stopping criterion based on a local linear trend diagnostic applied component-wise to the parameters optimization history. Formally, given a fixed window size $T\geq 2$ and any $t\geq T-1$, let
$\{\theta_{j}^{(s)}\}_{s=t-T+1}^{t}$
denote the last $T$ iterates of the $j$-th component. We consider the linear model with intercept
$\theta_{j}^{(s)}=a_{j}+c_{j}s+\varepsilon_{j}^{(s)},\qquad s\in\{t-T+1,\dots,t\},$
and compute the associated coefficient of determination $R_{j,t}^{2}$. Since in such a simple linear regression, the $R^{2}$ coefficient coincides with the squared correlation between the response and the regressor, small values of $R_{j,t}^{2}$ indicate that the recent evolution of $\theta_{j}$ is not better explained by a linear trend than by a constant model, and is therefore consistent with local stationarity.
The optimization procedure is stopped as soon as
$\frac{1}{\dim\theta^{(t)}}\sum_{j=1}^{\dim\theta^{(t)}}R_{j,t}^{2}\leq\delta,$
where $\delta\in[0,1]$ is a prescribed threshold (in practice $\delta=0.1$). With a relatively high value of $T=100$, this criterion that over the last $T$ iterations, no component displays a statistically meaningful linear drift, so the parameter vector is locally indistinguishable from a constant trajectory. In our findings, it proved to be robust in spite of the stochastic nature of the optimization path, and is scale-invariant.
To illustrate the convergence of the estimator, we examine the optimization process for a particular run in Figure 10, as well as RMSE values computed from $n_{\mathrm{runs}}=100$, as shown in Table 5.
Parameter True value Bias Std. deviation RMSE
$\gamma_{1}$ 2.000 -0.018 0.174 0.175
$\gamma_{2}$ 0.200 0.012 0.120 0.121
$\gamma_{3}$ 1.000 -0.009 0.101 0.101
$\frac{1}{2}\log(Q_{1})$ 0.949 -0.017 0.159 0.160
$\frac{1}{2}\log(Q_{2})$ 1.498 -0.035 0.237 0.240
$\frac{1}{2}\log(Q_{3})$ 1.151 -0.040 0.264 0.267
$\frac{1}{2}\log(R)$ 1.498 0.000 0.009 0.009
$\log\lambda^{1\to 1}$ -1.609 -0.009 0.152 0.152
$\log\lambda^{1\to 2}$ -4.605 -0.035 0.176 0.180
$\alpha_{1}^{1\to 1}$ -1.000 0.007 0.148 0.148
$\alpha_{1}^{1\to 2}$ 0.500 -0.010 0.144 0.144
$\beta^{1\to 1}$ -1.000 0.025 0.153 0.155
$\beta^{1\to 2}$ 0.500 0.018 0.143 0.144
Figure 5: Comparison of true and estimated parameters over $100$ runs with $n=500$ simulated individuals (three decimal places). Since the covariance matrix $Q$ is assumed diagonal, the $\log$-Cholesky reduces to the vector of negative $\log$-standard deviations.
The simulation results indicate that the proposed estimation procedure accurately recovers all parameters of the multi-state joint model for moderate sample sizes. Across the $100$ replications, the biases remain remarkably low, and the RMSE values are small compared to the magnitude of the corresponding parameters. Moreover, the observed variability is probably largely attributable to the intrinsic sampling variability of the true maximum likelihood estimator rather than to optimization error.
The population-level parameters $\gamma$ and the measurement noise parameter $R$ display comparatively lower dispersion, indicating stronger sensitivity for these components within the model. The optimization procedure exhibits stable convergence (Figure 10) and further supports that the interpretation that residual fluctuations across replications are primarily driven by the variance of the estimator itself.
6.3 Model Selection
To assess the performance of the information criteria described in Subsection 4.2, we compared the Akaike Information Criterion and the modified $\mathrm{BIC}\mathcal{I}$. Following the same procedure as in the rest of the current section, we performed $n_{\mathrm{runs}}=100$ independent simulations, generating for each run a dataset of size $n=500$ from the above specified model. For each dataset, we then fitted four competing models:
  1. 1.
    Correct model: the model used to generate the data, with correctly specified parameters.
  2. 2.
    Shared-$\alpha$ model: a simplified model in which the link parameters $\alpha^{k\to k^{\prime}}$ are constrained to be equal across all transitions, i.e., $\alpha^{1\to 1}=\alpha^{1\to 2}=\alpha$.
  3. 3.
    Expanded-$\beta$ model: a slightly more complex model in which three additional normally distributed covariates are added to each individual without any association.
  4. 4.
    Link misspecified: a model in which the link functions are incorrectly specified as the bi-exponential function instead of the cumulative exposure.
For each run, the model with the smallest value of AIC or $\mathrm{BIC}_{\mathcal{I}}$ was selected. Table 6 summarizes the results across the $100$ simulations.
Model AIC selections $\mathrm{BIC}_{\mathcal{I}}$ selections
Correct 69 99
Expanded-$\beta$ 31 1
Shared-$\alpha$ 0 0
Misspecified link 0 0
Figure 6: Number of times each model was selected based on AIC and $\mathrm{BIC}_{\mathcal{I}}$ over $100$ simulation runs.
The results indicate that both information criteria successfully identify the correct model in the majority of cases. However, $\mathrm{BIC}_{\mathcal{I}}$ favors the correctly specified model far more frequently than AIC, aligning with its stronger penalty for model complexity, and is able to correctly identify the model $99\%$ of the times.
Indeed, the increase in the AIC penalty for the expanded model is only $6$ relative to the correctly specified model. Yet the negative log-likelihood is of the order of 2 0002\,000. In such a regime, the additional penalty appears negligible compared with the scale of the likelihood term, so the AIC does not provide a sufficient criterion in this situation. Both the shared-$\alpha$ model and the misspecified link models were never selected by the AIC nor the $\mathrm{BIC}_{\mathcal{I}}$.
Overall, these findings demonstrate that our model selection criteria very effectively distinguish between correctly specified and misspecified multi-state joint models, providing reliable guidance for model choice in finite-sample settings.
7 Application to the PAQUID Cohort
The data used in this section originates from the PAQUID cohort [Letenneur1994], a large prospective population-based study initiated in southwestern France in 1988, aimed at understanding the determinants and trajectories of aging. A subsample of $n=500$ individuals was followed over a period of up to $20$ years [Proust-Lima2017], with repeated measures of cognitive and physical health, as well as socio-demographic characteristics.
In particular, global cognitive functioning was assessed through the Mini-Mental State Examination (MMSE), while physical dependency was evaluated using the HIER scale, which classifies subjects into four ordered states (see Figure 7): 0 (no dependency), $1$ (mild dependency), $2$ (moderate dependency), and $3$ (severe dependency). Individual covariates correspond to a pair of binary variables indicating whether or not the individual has a diploma and is a male $\left(\mathds{1}_{\mathrm{CEP}}(i),\mathds{1}_{\mathrm{male}}(i)\right)\in\{0,1\}^{2}$.
In this work, we focus on the association between cognitive decline and the progression of physical dependency by jointly modeling the longitudinal trajectory of MMSE scores and the transitions between the four HIER states using a joint multi-state model. More specifically, the state of an individual at a given time $t$ is defined as the highest HIER dependency score between trial entry and $t$, and progression is therefore monotone. This approach allows us to characterize the dynamic interrelationship between cognitive and physical aging processes within a unified statistical framework.
7.1 Preprocessing
Longitudinal values were normalized to the interval $[0,1]$ by dividing each score by its maximum value of $30$, and the measurement times were normalized to represent the elapsed time since age $65$ in decades. This technique ensured stable and fast convergence of the optimization process by keeping all parameters on roughly the same scale.
7.2 Model Specification
The transition graph was chosen as follows in Figure 7, and only allows monotonic transitions from a lower level of physical dependency to a higher one, as some transitions occur so infrequently that they cannot be meaningfully analyzed. We considered Weibull baseline hazards for all transitions, consistent with previous observations by Proust-Lima2017, with a clock-reset (sojourn time) specification. The parameters of these baseline hazards were jointly optimized with the model parameters. It should also be noted that not all patients begin in state 0, thus, in this respect, our model inherently accounts for left censoring.
Figure 7: Schematic representation of the four ordered HIER states, with possible transitions. HIER quantifies physical dependency (0: no dependency to 33: severe dependency) [Proust-Lima2017].
The regression and link functions were taken to be shifted and scaled logistic curve such that the normalized MMSE score is always is non-increasing
[…]
For every transition $(k,k^{\prime})\in E$, we define $g^{k\to k^{\prime}}:=g$. The parameter $\Psi_{1}$ governs the amplitude of the logistic function, $\Psi_{2}$ introduces a time shift, and $\Psi_{3}$ controls the slope at the inflection point.
Moreover, we imposed a specific substructure on the model by setting $\forall(k,k^{\prime})\in E,\,\beta^{k\to k^{\prime}}=\beta$, which helped mitigate issues arising from the small number of observations in some groups.
Additionally, since no prior information was available for the covariance matrix of the latent effects $Q\in\mathbb{R}^{3\times 3}$, we used a full $\log$-Cholesky parametrization of the precision matrix $Q^{-1}$, with six parameters denoted $\tilde{Q}_{1}$ through $\tilde{Q}_{6}$. The full parametrization is essential to capture the potential cross-dependencies between the three latent effects.
7.3 Results
In this study, the full dataset was used to estimate the inferential results presented in Subsection 7.3.1. For the dynamic prediction analysis in Subsection 7.3.2, we employed a 5-fold cross-validation procedure, corresponding to splits of $400$ observations for model training and $100$ for validation in each fold, as to obtain unbiased estimates of the predictive performance metrics.
7.3.1 Inference
The model was fitted on the complete dataset using the Adam optimizer and the optimization process was run until convergence based on the Equation 6.2 as in the simulation study. The Fisher Information Matrix was computed, implementing the method described in [baey2023efficient]. The standard errors were then computed as the square root of the diagonal of the inverse of the Fisher Information Matrix. Since, under the chosen parametrization, the value 0 lies in the interior of the admissible parameter space, Wald statistics can be used to test the nullity of each coefficient through the usual asymptotic normal approximation. However, we report only the significance results for the linear coefficients coefficients in either the covariates or the link functions, as they allow a direct interpretation. They are indicated in bold in the summarized Table 8. The optimization process itself is provided in the Appendix (see Figure 11).
Parameter Inferred value Std. error Parameter Inferred value Std. error
$\gamma_{1}$ 1.201 0.035 $\boldsymbol{\gamma_{2}}$ 0.221 0.042
$\boldsymbol{\gamma_{3}}$ 0.155 0.041 $\gamma_{4}$ 3.321 0.076
$\gamma_{5}$ 0.181 0.094 $\boldsymbol{\gamma_{6}}$ 0.568 0.119
$\gamma_{7}$ 0.636 0.188 $\boldsymbol{\gamma_{8}}$ 0.663 0.229
$\boldsymbol{\gamma_{9}}$ 1.168 0.233 $\tilde{Q}_{1}$ 2.537 0.079
$\tilde{Q}_{2}$ -0.032 0.321 $\tilde{Q}_{3}$ 0.644 0.086
$\tilde{Q}_{4}$ -1.909 0.134 $\tilde{Q}_{5}$ -0.752 0.100
$\tilde{Q}_{6}$ -0.690 0.065 $\frac{1}{2}\log(R)$ 3.023 0.008
$\log\lambda^{0\to 1}$ -1.174 1.090 $\log k^{0\to 1}$ 0.269 0.125
$\log\lambda^{0\to 2}$ 1.160 0.585 $\log k^{0\to 2}$ 0.648 0.142
$\log\lambda^{0\to 3}$ 1.663 0.769 $\log k^{0\to 3}$ 0.876 0.352
$\log\lambda^{1\to 2}$ 1.260 0.519 $\log k^{1\to 2}$ 0.434 0.097
$\log\lambda^{1\to 3}$ 2.744 0.926 $\log k^{1\to 3}$ 0.520 0.283
$\log\lambda^{2\to 3}$ 2.239 0.346 $\log k^{2\to 3}$ 0.509 0.107
$\alpha^{0\to 1}$ 1.254 1.614 $\boldsymbol{\alpha^{0\to 2}}$ -3.844 1.281
$\boldsymbol{\alpha^{0\to 3}}$ -8.487 1.362 $\boldsymbol{\alpha^{1\to 2}}$ -2.167 0.898
$\boldsymbol{\alpha^{1\to 3}}$ -8.393 0.957 $\boldsymbol{\alpha^{2\to 3}}$ -5.315 0.516
$\beta_{1}$ -0.159 0.315 $\beta_{2}$ 0.152 0.233
Figure 8: Inferred parameters and their estimated standard deviations. Bold values mean $p<0.05$ (only the linear coefficients included).
All link parameters $\alpha^{k\to k^{\prime}}$ for transitions $(k,k^{\prime})\in E$ are negative and statistically significant, except that from state 0 to $1$, and the most negative coefficients correspond to transitions reflecting rapid deterioration ($0\to 3$, $1\to 3$, and $2\to 3$). This pattern indicates that the current value of the MMSE score strongly influences the risk of progressing toward physical dependence, and lower cognitive performance is associated with a substantially increased transition intensity, especially for direct or fast declines.
Consistent with other analyses conducted on this dataset, both education level (CEP) and female gender are associated with more favorable cognitive and functional trajectories. Specifically, the inferred values of $\gamma_{2}$, $\gamma_{5}$, and $\gamma_{8}$ indicate that individuals with a primary school certificate (CEP) not only start from a significantly higher baseline level of MMSE, but also experience a delayed and slower progression of dementia. Similarly, the positive values of $\gamma_{3}$, $\gamma_{6}$, and $\gamma_{9}$ suggest the same remarks for women relative to men. Regarding the association coefficient $\beta$, the relatively large standard errors imply that the remaining effects are modest and that the link function already captures most of the relevant information.
7.3.2 Multi-State Dynamic Prediction
After fitting the model on each training fold, we performed multi-state dynamic prediction on the corresponding validation fold. For each validation fold $V_{k}$, with $k\in\llbracket 5\rrbracket$, predictions were computed using a single Monte Carlo integration scheme as described in Section 5. We assessed the performance of our model using two metrics: the Brier Skill Score (BSS), which is one of the most standard metric in this context, as well as the accuracy. More specifically, for each new individual $i$, longitudinal measurements and trajectories were truncated at various landmarks $t$, and predictions were made for future time points $u\geq t$ beyond each truncation. We considered the quantity $\chi_{u}^{*}(\mathcal{T}_{i}^{*})$ as defined by Example 1, which is estimated using the information available up to $t$, i.e., $\left(\mathcal{Y}_{i}(t),\mathcal{S}_{i}(t)\right)$.
Because the complete trajectories are not fully observed in the test sample and the quantity $\chi_{u}^{*}(\mathcal{T}_{i}^{*})$ is not available for $u>C_{i}$, predictions must account for censoring. For a given landmark time $t$, we obtain a collection of probability estimates
$\{\hat{\chi}_{i,u}(t)\}\subset\Delta\left(\{0,1,2,3\}\right),$
each computed from the information available up to $t$. The Brier Skill Score for the fold $V_{k}$ with $k\in\llbracket 5\rrbracket$ is then defined by
$\mathrm{BSS}_{k}(t,u)\coloneqq 1-\frac{\mathrm{BS}_{k}(t,u)}{\mathrm{BS}_{k}^{(0)}(u)},$
where
$\begin{cases}\mathrm{BS}_{k}(t,u)\coloneqq\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\left\|\hat{\chi}_{i,u\wedge C_{i}}(t)-\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right\|^{2},\\ \mathrm{BS}_{k}^{(0)}(u)\coloneqq\min_{p}\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\left\|p-\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right\|^{2}.\end{cases}$
Hence, values of $\mathrm{BSS}_{k}(t,u)$ close to $1$ indicate a substantial improvement over the optimal constant prediction.
Similarly, the accuracy at $(t,u)$ on the fold $V_{k}$ is defined as
$\mathrm{Acc}_{k}(t,u)\coloneqq\frac{1}{\#V_{k}}\sum_{i\in V_{k}}\mathds{1}\left\{\operatorname*{\arg\max}_{j\in\{0,1,2,3\}}\left(\hat{\chi}_{i,u\wedge C_{i}}(t)\right)_{j}=\operatorname*{\arg\max}_{j\in\{0,1,2,3\}}\left(\chi_{u\wedge C_{i}}^{*}(\mathcal{T}_{i}^{*})\right)_{j}\right\},$
which corresponds to the proportion of correctly identified states.
Figure 9 illustrates our dynamic predictions framework evaluated at three distinct landmark times using both metrics estimated from our cross-validation procedure. Both metrics are averaged across the $5$ folds.
Figure 9: Average values on the 55-fold cross-validation of the Brier Skill Score (top) and accuracy (bottom) at each time point for different prediction landmarks (dotted lines). Colored bands correspond to $\mathrm{mean}\pm 1.65\mathrm{sd}$.
Across the three prediction landmarks, the Brier Skill Score remains above $40\%$ at $10$ years beyond the prediction landmark. This indicates a very strong improvement over the best constant baseline model, despite the limited information available at the time of prediction.
Similarly, the classification accuracy consistently exceeds $60\%$, often higher, thus the model correctly identifies the majority of states. This level of performance persists even for long-term horizons extending more than $20$ years into the future, including during the early phases of the study.
As expected, predictions made immediately following each truncation are the most precise, while uncertainty naturally increases as the prediction horizon extends further. Both metrics also get substantially better as the prediction landmarks include more data. Both the Brier Skill Score (BSS) and classification accuracy remain robust over time, providing strong evidence of the reliability and stability of our predictive framework.
8 Conclusion
We presented a general likelihood-based framework for joint modeling of nonlinear longitudinal biomarkers and multi-state survival processes. The model unifies both components on arbitrary directed graphs, extending classical joint models to accommodate Markov and semi-Markov structures, recurrent transitions, and nonlinear mixed-effects submodels. This formulation provides a flexible and rigorous foundation for linking longitudinal biomarker trajectories to complex event histories within a single coherent probabilistic model.
Simulation studies confirmed robust parameter recovery, and an application to the PAQUID cohort highlighted the ability to capture the interplay between cognitive decline and physical dependency.
Since the model may potentially involve many transitions, and therefore a large number of parameters, a strategy of parameter sharing across transitions could help reduce the overall number of parameters. This is especially beneficial in settings with limited data for certain transitions, or when biological or clinical knowledge suggests similar effects across multiple transitions. Parameter sharing improves statistical efficiency, reduces overfitting, and facilitates model interpretability.
Future directions include a spline-based parametrization of baseline hazards to allow for nonparametric estimation. This approach is particularly useful when the true baseline hazard is expected to be complex or non-monotonic, or when parametric forms (such as exponential or Weibull) may be too restrictive and lead to model misspecification.
Another promising direction is the development of efficient visualization tools for multi-state trajectories and dynamic predictions. Interactive representations of individual event histories, transition probabilities, and longitudinal biomarker evolution would greatly facilitate model interpretation and communication of results to applied researchers.
Funding
This work was partially funded by the Stat4Plant project ANR-20-CE45-0012.
Acknowledgments
The authors would like to thank Jean-Benoist Leger for helpful discussions.
Conflict of interest
The authors declare that they have no conflicts of interest.
Data availability
The data analyzed in this study are publicly available via the R package timeROC (dataset Paquid).
Appendix A Multi-State Simulation
In nonlinear joint models with known design and known parameters parameters $\theta=(\gamma,Q,R,\alpha,\beta)$, the occurrence times of events conditionally on the latent variables $b_{i}$ and the covariates $X_{i}$ can be easily sampled using the inverse cumulative distribution function transform method. This may be achieved through bisection or other root-finding algorithms.
The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property 1 by considering one transition at a time, according to Algorithm 3. The algorithm is similar to Gillespie’s algorithm [gillespie2007stochastic], and can also include a survival condition $T_{i1}\geq t_{i,\mathrm{cond}}$ that uses the Chasles relation.
Here, we deliberately avoid imposing any conditioning in order to preserve maximal generality, since the algorithm is intended to be applicable beyond the specific framework of our model, only assuming continuous transition times distributions and the semi-Markov property.
Algorithm 3 Simulation of trajectory $i$
1:subject-specific censoring time $C_{i}\in\bar{\mathbb{R}}^{+}$; initial time-state pair $(T_{i0},S_{i0})$; optional survival condition $T_{i1}\geq t_{i,\mathrm{cond}}$, defaults to $t_{i,\mathrm{cond}}=T_{i0}$.
2:
3:Initialize $\ell\leftarrow 1$, $\mathcal{T}_{i}\leftarrow\left((T_{i0},S_{i0})\right)$
4:
5:while $T_{i(\ell-1)}<C_{i}$ and $\{j:(S_{i(\ell-1)},j)\in E\}\neq\emptyset$ do
6:  for each $j$ with $(S_{i(\ell-1)},j)\in E$ do
7:   Draw $T_{i\ell}^{(j)}$ such that $\forall t\geq T_{i(\ell-1)}\vee t_{i,\mathrm{cond}}$
8:      $\log\mathbb{P}(T_{i\ell}^{(j)}>t)=\int_{T_{i(\ell-1)}\vee t_{i,\mathrm{cond}}}^{t}\lambda_{i}^{S_{i(\ell-1)}\to j}(s\mid T_{i(\ell-1)})\,ds$
9:  end for
10:
11:  Set $T_{i\ell}\leftarrow\min_{j}T_{i\ell}^{(j)}$, and $S_{i\ell}\leftarrow\arg\min_{j}T_{i\ell}^{(j)}$
12:  Append $(T_{i\ell},S_{i\ell})$ to $\mathcal{T}_{i}$
13:  $\ell\leftarrow\ell+1$
14:end while
15:
16:if $T_{i(\ell-1)}>C_{i}$ then
17:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$
18:end if
19:return $\mathcal{T}_{i}$
The proof of the exactness of this algorithm is given below.
Proof.
First, consider the case $t_{i,\mathrm{cond}}=T_{i0}$. Let $\ell\geq$ and $t\geq T_{i\ell}$. One has
$\mathbb{P}(T_{i(\ell+1)}>t\mid T_{i\ell},S_{i\ell})$ $=\mathbb{P}\left(\bigcap_{j:(S_{i(\ell-1)},j)\in E}T_{i(\ell+1)}^{(j)}>t\mid T_{i\ell},S_{i\ell}\right),$
$\overset{\mathrel{\perp\!\!\!\perp}}{=}\prod_{j:(S_{i\ell},j)\in E}\mathbb{P}(T_{i(\ell+1)}^{(j)}>t\mid T_{i\ell},S_{i\ell}),$
$=\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i(\ell-1)}\to j}(T_{i\ell},t)\right).$
Furthermore, for any $k^{\prime}\in V$ such that $(S_{i\ell},k^{\prime})\in E$, calling $J_{i\ell}^{(k^{\prime})}\coloneqq\{j:(S_{i\ell},j)\in E\}\setminus\{k^{\prime}\}$, we have
$\mathbb{P}(S_{i(\ell+1)}=k^{\prime}\mid T_{i(\ell+1)}=t,T_{i\ell},S_{i\ell})$
$\propto p\left(T_{i(\ell+1)}^{(k^{\prime})}=t,\forall j\in J_{i\ell}^{(k^{\prime})}\,T_{i(\ell+1)}^{(j)}\geq t\mid T_{i\ell},S_{i\ell}\right)$
$=\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},t)\exp\left(-\sum_{j:(S_{i\ell},j)\in E}\Lambda_{i}^{S_{i\ell}\to j}(T_{i\ell},t)\right)$
$\propto\lambda_{i}^{S_{i\ell}\to k^{\prime}}(T_{i\ell},t).$
Combining both, we retrieve the same joint density as in Equation 2. If $t_{i,\mathrm{cond}}>T_{i0}$, one can also check that both densities are equal by simply integrating from $t_{\mathrm{cond}}$ instead of $T_{i\ell}$. Note that the condtioning only affects the first transition, as the next time $\forall\ell>0,\,T_{i\ell}\geq t_{i,\mathrm{cond}}\implies\forall\ell>0,\,T_{i\ell}\vee t_{i,\mathrm{cond}}=T_{i\ell}$. ∎
Appendix B Expression of the Semi-Markov likelihood
Proof.
The proof is short and relies on the fact that, using the semi-Markov property 1
$p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})_{\ell\leq m_{i}-1}\mid b_{i},X_{i}\right)$ $=\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell^{\prime}},S_{i\ell^{\prime}})_{\ell^{\prime}\leq\ell}\right),$
$=\prod_{\ell=0}^{m_{i}-1}p_{\theta}\left((T_{i(\ell+1)},S_{i(\ell+1)})\mid b_{i},X_{i},(T_{i\ell},S_{i\ell})\right).$
On the other hand, the probability that no additional event is observed between $T_{im_{i}}$ and $C_{i}$ is $1$ if the last state reached is absorbing, and otherwise, since censoring is assumed non-informative, it can be treated as an independent constant
$\mathbb{P}_{\theta}\left(T_{i(m_{i}+1)}>C_{i}\mid b_{i},X_{i},(T_{im_{i}},S_{im_{i}})\right)=\exp\left(-\sum_{j:(S_{im_{i}},j)\in E}\Lambda_{i,\theta}^{S_{im_{i}}\to j}(T_{im_{i}},C_{i}\mid b_{i},X_{i})\right),$
so we recover the formula cited above, with the convention that an empty sum is 0. ∎
Appendix C Stochastic optimization process on simulated data
Figure 10: Evolution of the parameters during the optimization of the marginal $\log$-likelihood using stochastic gradient ascent. Dotted horizontal lines correspond to the true values, while vertical lines mark the start of each window used to evaluate the stopping criterion.
Appendix D Stochastic optimization process on the PAQUID dataset
Figure 11: Evolution of the parameters during the optimization of the marginal $\log$-likelihood using stochastic gradient ascent. Vertical lines mark the start of each window used to evaluate the stopping criterion.